### v1

# Implementing Cumulative Functions with Generalized Cumulative Constraints

Abstract
Modeling scheduling problems with conditional time intervals and cumulative functions has become a common approach when using modern commercial constraint programming solvers. This paradigm enables the modeling of a wide range of scheduling problems, including those involving producers and consumers. However, it is unavailable in existing open-source solvers and practical implementation details remain undocumented. In this work, we present an implementation of this modeling approach using a single, generic global constraint called the Generalized Cumulative. We also introduce a novel time-table filtering algorithm designed to handle tasks defined on conditional time-intervals. Experimental results demonstrate that this approach, combined with the new filtering algorithm, performs competitively with existing solvers enabling the modeling of producer and consumer scheduling problems and effectively scales to large problems.
Introduction
The success of Constraint Programming as a modeling and solving technology for hard scheduling problems is well established (Laborie et al. 2018). The modeling has been further enhanced with the introduction of conditional time-interval variables (Laborie and Rogerie 2008), which represent the execution of tasks that may or may not be executed. Such variables allow to conveniently model problems with alternative resources or optional tasks such as in (Kinable, Wauters, and Berghe 2014; Kumar et al. 2018; Cappart and Schaus 2017; Kizilay, Eliiyi, and Van Hentenryck 2018). Building on conditional time-intervals, the same authors introduced an algebraical model for cumulative resources called cumulative functions (Laborie et al. 2009). It provides a convenient way to model producer-consumer problems with optional tasks such as in (Liu, Aleman, and Beck 2018; Gedik et al. 2018; Cappart et al. 2018; Thomas and Schaus 2024). CPOptimizer (Laborie et al. 2018) and OptalCP (Vilím and Pons 2023) are the only two solvers that fully support conditional time-interval variables and cumulative functions and this modeling paradigm is not supported by existing open-source solvers [^1] API languages such as Minizinc also support optional variables at the modeling layer (Mears et al. 2014).
We propose an implementation of this modeling paradigm with lower and upper-bound restrictions on cumulative functions using a generalized cumulative constraint that supports negative heights (Beldiceanu and Carlsson 2002). This more general form of the cumulative constraint allows variable and negative resource consumptions as well as optional tasks. We introduce a new filtering algorithm for this constraint that is based on time-tabling (Gay, Hartert, and Schaus 2015a; Fahimi, Ouellet, and Quimper 2018; Letort, Carlsson, and Beldiceanu 2015) and does additional filtering on the height and length of tasks. The algorithm proceeds in two steps. First, it uses the Profile data structure recently introduced by (Gingras and Quimper 2016) and builds the optimistic/pessimistic profiles of resources produced/consumed by tasks. It then processes each task to prune its bounds by comparing it to the profile. The experimental results show that the newly introduced algorithm performs well compared to existing ones, particularly in large-scale scheduling problems solved with a greedy search.
Modeling with Cumulative Functions
Conditional Time-Intervals
The domain of a conditional time-interval variable can be expressed as a subset of {}$\{\bot\}\cup\{[s,e)\mid s\leq e,s,e\in\mathbb{Z}\}$ (Laborie and Rogerie 2008). It is fixed if $x=\bot$ (the interval is not executed) or $x=[s,e)$ (the interval starts at $s$ and ends at $e$). For convenience, the description of the filtering algorithm manipulates conditional time-interval domains through the following attributes for each task $i\in T$:
  • $p_{i}\in\{true,false\}$ is the execution status ($false=\bot$).
  • $s_{i}\in[\underline{s}_{i},\overline{s}_{i}]$ and $e_{i}\in[\underline{e}_{i},\overline{e}_{i}]$ are the start/end variables.
  • $d_{i}\in[\underline{d}_{i},\overline{d}_{i}]$ is the (non negative) duration variable.
A bound consistent filtering on the relation $s_{i}+d_{i}=e_{i}$ is maintained atomically on the corresponding attributes, and the domain of a time-interval variable becomes $\bot$ whenever one attribute range ($s,d$ or $e$) becomes empty which may trigger an inconsistency if $p_{i}=true$. The whole domain of a conditional time-interval $i\in T$ is thus encoded by a tuple of variables $x_{i}=\langle s_{i},d_{i},e_{i},p_{i}\rangle$.
Cumulative Functions
For modeling cumulative problems, (Laborie et al. 2009) introduced the idea of cumulative function expressions. They allow the expression of step-wise integer functions of conditional time-interval variables. The contribution of a single time-interval variable $i\in T$ is represented by an elementary cumulative function receiving an interval $x_{i}$ and a non-negative consumption value $c_{i}$ or a non-negative height range $[\underline{c}_{i},\overline{c}_{i}]$. The two main [^2] Variants exist such as stepAtEnd. elementary functions are pulse and stepAtStart (Laborie et al. 2009) as represented in Figure 1.
Figure 1: Elementary cumulative functions.
Cumulative functions (elementary or not) can then be combined with the plus/minus operators to form a more complex function representing the addition or subtraction of two functions. A cumulative function can thus be viewed as an Abstract Syntax Tree (AST) with elementary functions at the leaf nodes.
Example 1.
Three tasks $A$, $B$ and $C$ contribute to a cumulative function $f=stepAtStart(A,2)\allowbreak-\allowbreak(pulse(B,1)\allowbreak+\allowbreak stepAtEnd(C,1))$. The resulting AST and cumulative function are represented in Figure 2. On the right, the solid parts represent the time windows of the tasks while the hatched parts represent their contribution to the profile. A dotted line represents the final cumulative function $f$.
Figure 2: Example 1: AST (left) and cumul. function (right).
A cumulative function can be flattened using Algorithm 1 to extract a set of cumulative tasks (tasks with their positive or negative heights). This algorithm traverses the AST and collects the activities at the leaf nodes, reversing the height signs for the activities flattened on the right branch of a minus node. Note that whenever encountering a $stepAtStart(x,c)$ leaf node, the cumulative task collected is defined on a fresh task interval $x^{\prime}$ related to the original one with: $x^{\prime}.s=x.s$ (same start as $x$), $x^{\prime}.p=x.p$ (same execution status), and $x^{\prime}.e=Horizon$ (ending at the time horizon) to reflect the definition of a step function [^3] Other cases such as $stepAtEnd(x,c)$ are defined similarly..
1 if $n=\text{Pulse}(x,c)$ then  $\{(x,c)\}$
2 else if $n=\text{StepAtStart}(x,c)$ then  $\{(x^{\prime},c)\}$
3 else if $n=\text{Plus}(n_{l},n_{r})$ then
4 $\text{flatten}(n_{l})\cup\text{flatten}(n_{r})$
5else if $n=\text{Minus}(n_{l},n_{r})$ then
6 $\text{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\text{flatten}(n_{r})\}$
7else  $error$ // Unknown pattern
[…]
Example 2.
Continuing example 1, $f$) collects the cumulative tasks $\{(A^{\prime},2),(B,-1),(C^{\prime},-1)\}$ illustrated on the right of Figure 2.
Constraints on Cumulative Functions
The height of a cumulative function can be restricted with a lower and upper bound at every time overlapping at least one executing time-interval: $\underline{C}\leq f\leq\overline{C}$. This constraint is called $alwaysIn(f,\underline{C},\overline{C})$ in (Laborie et al. 2009). It can be enforced by posting a GeneralizedCumulative constraint requiring a resource consumption between $\underline{C}$ and $\overline{C}$ for the set of cumulative tasks $T$ collected with the Flatten function. The whole domain of a conditional (cumulative) task $i\in T$ is encoded by a tuple of variables $\langle s_{i},d_{i},e_{i},c_{i},p_{i}\rangle$ where the integer variable $c_{i}$ represents its consumption (possibly negative or hybrid) on the resource. The set of tasks $T$ is partitioned into three subsets as given in Definition 1.
Definition 1.
The set $O=\{i\in T\mid p_{i}=\{true,false\}\}$ denotes the set of optional tasks, those for which it is not decided yet if they will execute or not on the resource. The set $R=\{i\in T\mid p_{i}=true\}$ denotes the set of required tasks, those which will surely be executed on the resource. The set $E=\{i\in T\mid p_{i}=false\}$ denotes the set of excluded tasks, those which will surely not be executed on the resource. The sets $O$, $R$ and $E$ are disjoint and form a partition of the set $T$, i.e., $T=O\cup R\cup E$.
Definition 2.
[…]
$\sum\limits_{i\in R\mid s_{i}\leq\tau<e_{i}}c_{i}\in[\underline{C},\overline{C}]\qquad\forall\tau\in\bigcup_{i\in R}[s_{i},e_{i})$ (1)
Restricted Time Interval Scope
To enforce the capacity range at every time point across the entire horizon rather than only when at least one task executes, one can add a dummy task that spans the entire horizon, with a mandatory execution status and a height of zero. More generally, the time scope can also be limited to a specific task, by linking the status variables of the tasks to the overlap of this task.
Consistency
Enforcing global bound-consistency for the cumulative constraint is NP-Hard (Garey and Johnson 1979). However, enforcing it on the decomposition into sum constraints at every point in the horizon, as in (1), is not NP-Hard. Unfortunately, this approach is inefficient for large horizons, as it requires a sum constraint to be enforced at every time point. Therefore, we aim to achieve the same filtering in polynomial time relative to the number of tasks, rather than the size of the horizon. This filtering is generally referred to as time-tabling filtering. Intuitively, it works by aggregating all the fixed parts (when $\underline{s}_{i}<\overline{e}_{i}$) to postpone task execution.
Timetabling Algorithm
The Profile Data Structure
The Profile data structure is an aggregation of adjoined rectangles of different lengths and heights introduced in (Gingras and Quimper 2016; Gay, Hartert, and Schaus 2015a) to record the resource utilization of tasks. The end of a rectangle is the beginning of the next one. The starting (resp. ending) of a rectangle is represented by a tuple called time point with a component representing the beginning (resp. ending) time of the rectangle and other components that contain information relative to the resource consumption for the duration of the rectangle. These time points are sorted in increasing order of time and are kept in a linked list structure called Profile. An original idea of (Gingras and Quimper 2016) is to have pointers $tp(\underline{s}_{i})$, $tp(\underline{e}_{i})$ and $tp(\overline{e}_{i})$ linking to the time point associated with $\underline{s}_{i}$, $\underline{e}_{i}$, and $\overline{e}_{i}$ respectively.
Timetable Filtering
[…]
Intuitively, the minimum profile is obtained by considering the upper bound of the negative contribution and the lower bound of the positive contribution of each task. In contrast, the maximum profile considers the upper bound of the positive contribution and the lower bound of the negative contribution of each task. The upper bound of the negative (resp. positive) contribution of a task corresponds to its largest negative (resp. positive) height during the whole possible time window of the task while the lower bound of its negative (resp. positive) contribution corresponds to its smallest negative (resp. positive) height during the fixed part of the task. More formally, the minimum $\underline{P}$ (resp. maximum $\overline{P}$) resource profile at time $t$ is defined as
$\underline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\min(\underline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\max(\underline{c}_{i},0)$ (2)
$\overline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\max(\overline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\min(\overline{c}_{i},0)$ (3)
where $i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}$ corresponds to the set of all optional and required tasks whose time window overlaps the time $t$ and $i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}$ corresponds to the set of all required tasks that have a fixed part that overlaps the time $t$.
The proposed timetable filtering algorithm is split into two steps: First, the profile data structure $P$ is initialized by creating and ordering the time points, linking them with the tasks, computing their minimum and maximum profile ranges, and checking their consistency by comparing the profile ranges with the maximum and minimum capacity. Second, each task is checked during its entire possible span in terms of profile and capacity to adjust its domain.
Timeline Initialization and Consistency Check
The function initializeTimeline (Algorithm 2) initializes the timeline, links the corresponding time points to tasks, and computes the minimum and maximum profile of the resource utilization. It receives as input the set of tasks $T$. First, events are generated by iterating over the tasks that are not absent (Line 2). For each non-absent task $i$, two events are generated which correspond to the minimum start time ($\underline{s}_{i}$), and the maximum end time ($\overline{e}_{i}$) of the task. For tasks that are present and with a fixed part, two additional events are generated for the maximum start time ($\overline{s}_{i}$) and the minimum end time ($\underline{e}_{i}$) of the task.
1 ${}events\leftarrow\{\}$
2 for $i\in T\mid i\in O\cup R$ do
3 $events\leftarrow events\cup\{\underline{s}_{i},\overline{e}_{i}\}$
4 if $i\in R\wedge\overline{s}_{i}<\underline{e}_{i}$ then
5    $events\leftarrow events\cup\{\overline{s}_{i},\underline{e}_{i}\}$
[…]
7 $tp\leftarrow new\,tp(\min(events.time))$
8 $P\leftarrow\{tp\}$
9 for $e\in events$ sorted by time do
10 $i\leftarrow e.task$
11 if $tp.time<e.time$ then
12    if #$\#fp>0\wedge(\underline{P}>\overline{C}\vee\overline{P}<\underline{C})$ then
13       return Fail
14
15    $tp\leftarrow new\,tp(e.time)$
16    $P.append(tp)$
17
18 if $e=\underline{s}_{i}$ then
19    $\underline{P}\leftarrow\underline{P}+\min(\underline{c}_{i},0)$
20    $\overline{P}\leftarrow\overline{P}+\max(\overline{c}_{i},0)$
21    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
22    $tp(\underline{s}_{i})\leftarrow tp$
23
24 if $e=\overline{e}_{i}$ then
25    $\underline{P}\leftarrow\underline{P}-\min(\underline{c}_{i},0)$
26    $\overline{P}\leftarrow\overline{P}-\max(\overline{c}_{i},0)$
27    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
28    $tp(\overline{e}_{i})\leftarrow tp$
29
30 if $e=\overline{s}_{i}$ then
31    $\underline{P}\leftarrow\underline{P}+\max(\underline{c}_{i},0)$
32    $\overline{P}\leftarrow\overline{P}+\min(\overline{c}_{i},0)$
33    #$\#fp\leftarrow\#fp+1$
34    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
35
36 if $e=\underline{e}_{i}$ then
37    $\underline{P}\leftarrow\underline{P}-\max(\underline{c}_{i},0)$
38    $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
39    #$\#fp\leftarrow\#fp-1$
40    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
41
return $P$
[…]
Then, these events are processed by order of increasing time to initialize the timeline $P$ (Line 2). A current time point $tp$ is maintained and corresponds to the last time point of the timeline being initialized. New time points are created only when the time of the event differs from the current time point $tp$ as checked at Line 2. This means that if several events share the same time, they will all contribute towards the same time point. When a new time point is created, it is appended at the end of the timeline which links it to the previous last time point. When processing a start min ($\underline{s}$) or end max ($\overline{e}$) event, the current time point is linked to the event at Lines 2 and 2 to retrieve the time points corresponding to the start and end of tasks in constant time later.
[…]
As the timeline is initialized, the consistency of the minimum and maximum profiles is checked for each time point at Line 2. Once a time point is completed (when creating the next time point), we ensure that if at least one task is guaranteed to execute during the time point (#$\#fp>0$), the minimum profile $\underline{P}$ is not above the maximum capacity of the resource $\overline{C}$ and the maximum profile $\overline{P}$ is not below the minimum capacity of the resource $\underline{C}$.
The worst-case complexity in time of the function initializeTimeline is $\mathcal{O}(n\log(n))$ and is due to the sorting of the events used in the loop of Line 2. Example 3 illustrates the profile initialization for three tasks.
Example 3.
Let us consider three tasks :
$A$
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=[1,2],\allowbreak p=true\rangle$;
$B$
: $\langle s=[2,4],\allowbreak d=[3,4],\allowbreak e=[5,7],\allowbreak c=2,\allowbreak p=true\rangle$;
$C$
: $\langle s=[3,8],\allowbreak d=[1,3],\allowbreak e=[4,9],\allowbreak c=[-2,1],\allowbreak p=\{true,false\}\rangle$;
and a resource $C$ of capacity bounds [0,1][0,1][ 0 , 1 ]. The time points and their attributes are shown in Table 1. Figure 3 shows a representation of the tasks and the resulting profile ranges. Note that the events $\overline{s}_{c}$ and $\underline{e}_{c}$ are not considered as task $C$ has no fixed part.
[…]
0 $\underline{s}_{a}$ 0 2 0
1 $\overline{s}_{a}$ 1 2 1
2 $\underline{s}_{b}$ 1 4 1
3 $\underline{e}_{a},\underline{s}_{c}$2 5 0
4 $\overline{e}_{a},\overline{s}_{b}$ 0 3 1
5 $\underline{e}_{b}$2 3 0
7 $\overline{e}_{b}$2 1 0
9 $\overline{e}_{c}$ 0 0 0
Table 1: Time points computed from the tasks of Example 3.
Figure 3: Representation of the tasks of Example 3 with the resulting minimum ($\underline{P}$) and maximum ($\overline{P}$) profile.
Filtering
Once the timeline is initialized and its consistency verified, the profile range (minimum and maximum profile of tasks at any time point) is used to filter the time windows, capacity, and duration of tasks. Four different filtering rules are used:
Forbid
The first rule is used to reduce the time windows of tasks by checking for each task $i\in T$ if starting (resp. ending) the task at its minimum start time (resp. maximum end time) is possible in regards to the profile range and the capacity range. If this is not the case, the task is pushed to the earliest (resp. latest) time at which it can be placed without violating the capacity range. Formally, the rule is defined as:
$\forall tp\in P\,|\,\underline{s}_{i}\leq tp.time<\min(\overline{s}_{i},\underline{e}_{i}),$
$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (4)
$\Rightarrow\underline{s}_{i}\geq tp.next.time$
$\forall tp\in P\,|\,\max(\overline{s}_{i},\underline{e}_{i})\leq tp.time<\overline{e}_{i},$
$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (5)
$\Rightarrow\overline{e}_{i}\leq tp.prev.time$
Note that as tasks are modeled with interval variables, the rule $e_{i}=s_{i}+d_{i}$ is enforced internally in the variable. Thus, when a task $i$ has its minimum start time $\underline{s}_{i}$ (resp. maximum end time $\overline{e}_{i}$) adjusted, its minimum end time $\underline{e}_{i}$ (resp. maximum start time $\overline{s}_{i}$) is also updated by the relation $\underline{e}_{i}\geq\underline{s}_{i}+\underline{d}_{i}$ (resp. $\overline{s}_{i}\leq\overline{e}_{i}-\underline{d}_{i}$) embedded in the interval variable.
Mandatory
The second rule detects if the execution of a task $i\in T$ at a time point $tp$ is necessary to avoid a violation of the resource capacity range at this time point. If this is the case, the task is adjusted to make sure it is present during the time point: 1) if the task is optional it is set to present; 2) its time window is adjusted such that it executes over the whole time point; 3) its height is adjusted to avoid the profile range violation. In formal terms, the rule is defined as:
$\forall tp\in P\,|\,tp.time\in[\underline{s}_{i},\overline{e}_{i})\wedge tp.\#fp>0,$
$tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C}$
$\Rightarrow\begin{cases}&p_{i}\leftarrow true\\ &\overline{s}_{i}\leq tp.time\\ &\underline{e}_{i}\geq tp.next.time\\ &\underline{c}_{i}\geq deficit\text{ if }deficit>0\\ &\overline{c}_{i}\leq overload\text{ if }overload<0\end{cases}$ (6)
where $deficit=\underline{C}-(\overline{P}-\max(\overline{c}_{i},0))$ and $overload=\overline{C}-(\underline{P}-\min(\underline{c}_{i},0))$ are the deficit and overload of height needed for the profile range to intersect the capacity range.
Height
This rule adjusts the height of each task $i\in T$ in regard to the resource capacity range. The maximum (resp. minimum) value of the height of a task is bounded by the difference between the maximum (resp. minimum) capacity of the resource and the minimum (resp. maximum) profile without the contribution of the task. Two cases are possible:
1). If the task has a fixed part, its height is checked and adjusted directly at each time point of the fixed part:
$\forall tp\in P\mid\overline{s}_{i}\leq tp.time<\underline{e}_{i},$
$\overline{c}_{i}\leq\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$ (7)
$\underline{c}_{i}\geq\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
Note that if a task is present ($i\in R$), its minimum positive ($\max(\underline{c}_{i},0)$) and negative ($\min(\overline{c}_{i},0)$) contributions have been used to compute the profile range and thus these values must be subtracted in the above equation.
2). If the task has no fixed part, the maximum available height over the minimum overlapping interval of the task is used to adjust its height:
$\overline{c}_{i}\leq\max_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$
$\underline{c}_{i}\geq\min_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$ (8)
The minimum overlapping interval was introduced by (Gay, Hartert, and Schaus 2015b) and is defined for a task $i\in T$ as the interval $[\underline{e}_{i}-1,\overline{s}_{i}]$. Intuitively, it corresponds to the smallest interval such that, no matter its start time, length or end time, the task executes during at least one time unit of this interval.
Length
The last rule is inspired by (Ouellet and Quimper 2019). It adjusts the maximum duration $\overline{d}$ of the tasks that do not have a fixed part (for the tasks with a fixed part, the adjustment would be redundant with the Forbid rule). The principle is to find the longest time span during which the task can be scheduled without a resource capacity violation.
The function timetabling of Algorithm 3 checks these rules for each task $i\in T$. This algorithm receives as input the set of tasks $T$ and the bounds of the resource capacity $[\underline{C},\overline{C}]$. The algorithm iterates over all non-fixed tasks at Line 3.
1 $P\leftarrow\text{initializeTimeLine}(T)$
2 $O\leftarrow\{i\in T\mid p_{i}=\{true,false\}\}$
3 $R\leftarrow\{i\in T\mid p_{i}=true\}$
4 for $i\in O\cup R$ do
5 if $i$ is fixed then continue
6 $tpf\leftarrow tp(\underline{s}_{i})$
7 while $tpf.time<\min(\overline{s}_{i},\underline{e}_{i})$ do
8    if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
9       $\underline{s}_{i}\leftarrow\max(tpf.next.time,\underline{s}_{i})$
10
11    else
[…]
13
14    $tpf\leftarrow tpf.next$
15
16 $tpb\leftarrow tp(\overline{e}_{i}).prev$
17 while $tpb.next.time>\max(\overline{s}_{i},\underline{e}_{i})$ do
18    if $tpb.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpb.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
19       $\overline{e}_{i}\leftarrow\min(tpb.time,\overline{e}_{i})$
20
21    else
[…]
23
24    $tpb\leftarrow tpb.prev$
25
26 if $\overline{s}_{i}<\underline{e}_{i}$ then
27    while $tpf.time<\underline{e}_{i}$ do
[…]
29       $\underline{c}^{\prime}\leftarrow\underline{C}-(tpf.\overline{P}-\max(\overline{c}_{i},0))$
30       $\overline{c}^{\prime}\leftarrow\overline{C}-(tpf.\underline{P}-\min(\underline{c}_{i},0))$
31       if $i\in R$ then
32          $\underline{c}^{\prime}\leftarrow\underline{c}^{\prime}+\min(\overline{c}_{i},0)$
33          $\overline{c}^{\prime}\leftarrow\overline{c}^{\prime}+\max(\underline{c}_{i},0)$
34
35       $\underline{c}_{i}\leftarrow\max(\underline{c}^{\prime},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{\prime},\overline{c}_{i})$
36       $tpf\leftarrow tpf.next$
37
38 else
39    $\overline{d}^{*}\leftarrow 0$; $s^{\prime}\leftarrow\underline{s}_{i}$
40    $\underline{c}^{*}\leftarrow\underline{C}-tpf.prev.\overline{P}+\max(\overline{c}_{i},0)$
41    $\overline{c}^{*}\leftarrow\overline{C}-tpf.prev.\underline{P}+\min(\underline{c}_{i},0)$
42    while $tpf.time<\overline{s}_{i}$ do
43       $\overline{d}^{*}\leftarrow\max(tpf.time-s^{\prime},\overline{d}^{*})$
44       if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
45          $s^{\prime}\leftarrow tpf.next.time$
46
47       else
[…]
49
50       $\underline{c}^{*}\leftarrow\min(\underline{C}-tpf.\overline{P}+\max(\overline{c}_{i},0),\underline{c}^{*})$
51       $\overline{c}^{*}\leftarrow\max(\overline{C}-tpf.\underline{P}+\min(\underline{c}_{i},0),\overline{c}^{*})$
52       $tpf\leftarrow tpf.next$
53
54    $\overline{d}^{*}\leftarrow\max(\overline{e}_{i}-s^{\prime},\overline{d}^{*})$; $\overline{d}_{i}\leftarrow\min(\overline{d}^{*},\overline{d}_{i})$
55    $\underline{c}_{i}\leftarrow\max(\underline{c}^{*},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{*},\overline{c}_{i})$
56
[…]
Each task is processed in three steps:
First, the Forbid rule is used to adjust the minimum start time of the task by iterating forward over the time points in the interval $[\underline{s}_{i},\min(\overline{s}_{i},\underline{e}_{i}))$ in the loop at Line 3. Second, the Forbid rule is used to adjust the maximum end time of the task by iterating backward over the interval $[\max(\overline{s}_{i},\underline{e}_{i}),\overline{e}_{i})$ in the loop at Line 3.
The last step iterates over the time points of the remaining part of the time window. It depends if the task has a fixed part or not. In the former case, the loop at Line 3 iterates over the time points in the fixed part of the task and adjusts its height following the 1st case of the Height rule. In the case where the task has no fixed part, the loop at Line 3 is executed. The height of the task is adjusted according to the 2nd case of the Height rule and its maximum length is adjusted following the Length rule.
In order to compute the maximum length of the task, we maintain the variables $\overline{d}^{*}$ which contains the maximum length found so far and $s^{\prime}$ which corresponds to the earliest time at which the task can start and span without capacity violation until the current time point. During the loop, $\overline{d}^{*}$ is updated if the difference between the current time point time and $s^{\prime}$ is greater than its previous value. If the task is detected as infeasible during a time point, $s^{\prime}$ is set to the end of this time point (Line 3). Note that in order to take into account the parts of the time window that are not considered in Loop 3, $s^{\prime}$ is initialized to $\underline{s}$ and the difference between $\overline{e}$ and $s^{\prime}$ is considered at the end of the loop (Line 3).
Similarly, the minimum and maximum available heights are computed as $\underline{c}^{*}$ and $\overline{c}^{*}$ and used to update the height of the task if it has no fixed part (Line 3). Note that these variables are initialized based on the height available at the time point before the the start of Loop 3 in order to consider the whole minimum overlapping interval of the task.
The Mandatory rule is enforced over the whole time window of each task. The function checkIfMandatory of Algorithm 4 is used to check this rule and apply its adjustments if needed. It is called in each loop, at Lines 3, 3, 3 and 3.
1 if  #$\#fp>0\wedge(tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C})$  then
2 $p_{i}\leftarrow true$
3 $\overline{s}_{i}\leftarrow\min(tp.time,\overline{s}_{i})$
4 $\underline{e}_{i}\leftarrow\max(tp.next.time,\underline{e}_{i})$
5 $deficit\leftarrow\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
6 $overload\leftarrow\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
7 if $deficit>0$ then $\underline{c}_{i}\leftarrow\max(deficit,\underline{c}_{i})$
8 if $overload<0$ then $\overline{c}_{i}\leftarrow\min(overload,\overline{c}_{i})$
9
[…]
The worst-case time complexity of Algorithm 3 is $\mathcal{O}(n^{2})$. Note that some optimizations are possible depending on the set of tasks given as input. Indeed, the Mandatory propagation rule is only useful if the tasks have a mix of positive and negative heights or if there is a minimum capacity $\underline{C}$ to enforce. If all the height variables are fixed to a single value, the Height adjustment rule is not necessary. Similarly, the Length adjustement rule is only needed if the tasks have variable lengths. If these three rules can be ignored, only steps 1 and 2 need to be performed and the unnecessary loops at Lines 3 and 3 can be avoided.
Another optimization is to avoid considering some tasks when performing the timetabling algorithm. Indeed, fixed tasks that occur before the start of the earliest unfixed task or after the end of the latest unfixed task are not useful for filtering. Thus, such tasks do not need to be processed which reduces the number of time points in the profile. To do so, we use the same process as the Fruitless Fixed Tasks Removal of (Gay, Hartert, and Schaus 2015a).
Example 4.
When executing the timetabling algorithm on the same tasks as in Example 3, the adjustments are:
  1. 1.
    When processing task $A$ at time point 111, its maximum height $\overline{c}_{A}$ is adjusted to 111 by the Height rule.
  2. 2.
    When processing task $B$ at time point 222, the task is detected as non-feasible at this time point by the Forbid rule. Its minimum start $\underline{s}_{B}$ is thus pushed to the next time point at time 333 where the adjustment stops as the task is feasible at this time. This also updates its minimum end to $\underline{e}_{B}=6$ due to the internal relation $\underline{e}_{B}\geq\underline{s}_{B}+\underline{d}_{B}$.
  3. 3.
    When processing task $C$ at time point 444, the task is detected as mandatory. The following adjustments are done: 1) The task is set to present. 2) Its maximum starting time is set to $\overline{s}_{C}=4$ which also updates its maximum ending time to $\overline{e}_{C}=7$ due to the internal relation $\overline{e}_{C}\leq\overline{s}_{C}+\overline{d}_{C}$. 3) Its minimum ending time is set to $\underline{e}_{C}=5$. 4) As the value $overload=-1$ is negative, the maximum height of the task is set to $\overline{c}_{C}=-1$. Note that even if task $B$ is processed before task $C$ and its fixed part increased, task $C$ is adjusted based on the profile range which has been computed based on the state of task $B$ at the start of the propagation.
The state of the tasks after timetabling is:
$A$
: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=1,\allowbreak p=true\rangle$
$B$
: $\langle s=[3,4],\allowbreak d=[3,4],\allowbreak e=[6,7],\allowbreak c=2,\allowbreak p=true\rangle$
$C$
: $\langle s=[3,4],\allowbreak d=[1,3],\allowbreak e=[5,7],\allowbreak c=[-2,-1],\allowbreak p=true\rangle$
Related Work
In (Beldiceanu and Carlsson 2002), a timetabling filtering for the GeneralizedCumulative constraint with alternative resources was introduced. It collects a list of tasks at relevant time-point before filtering them. Our method is more in line with the strategy described in (Gay, Hartert, and Schaus 2015a) for the classical cumulative, which also constructs a profile and then confronts each task with it. We instead use the Profile data structure (Gingras and Quimper 2016), to prune the tasks forward and backward. This reduces the number of calls required to reach the fix-point of the constraint. Contrary to (Gay, Hartert, and Schaus 2015a), which uses just a list of rectangles for the profile, our algorithm finds the starting rectangle in $\mathcal{O}(1)$ thanks to the Profile. Consequently, despite a worst-case quadratic time complexity per call similar to that described in (Beldiceanu and Carlsson 2002) and (Gay, Hartert, and Schaus 2015a), the proposed algorithm can be faster. Furthermore, our algorithm also prunes the height and the maximum length of tasks. More details on the filtering differences with the other implementations from the literature are provided in the technical Appendix.
Experimental Results
Our approach and filtering algorithm was implemented in Java on MaxiCP, an extension of the MiniCP (Michel, Schaus, and Van Hentenryck 2021) solver. We compare it with the closed-source state-of-the-art commercial solver CPOptimizer (Laborie et al. 2018) as well as the open-source solver Gecode (Schulte, Lagerkvist, and Tack 2006) that implements a version of the GeneralizedCumulative constraint of (Beldiceanu and Carlsson 2002). As cumulative functions are not available in Gecode, we carefully modeled them in a way similar to how they would be flattened by Algorithm 1. Three cumulative problems that involve reservoirs or negative consumptions are considered. We use the same fixed static search strategy for all solvers and problems to ensure that only the filtering strength and propagation speed influence the results. The experiments were carried out on a MacBook Pro M3 with 32GB of memory. A description of the models is available in the technical appendix, and the sourcecodes are available in the supplementary material. Figure 4 shows the cumulative number of instances solved within a virtual time limit for the three problems. The appendix also presents a pairwise comparison of solvers on the number of backtracks for instances that were commonly solved.
RCPSP-CPR
The Resource Constrained Project Scheduling Problem with Consumption and Production of Resources (RCPSP-CPR) is a variant of the RCPSP with additional reservoir resources that must always be kept positive (Koné et al. 2013). Each task is present and, in addition to classical renewable resource consumption (modeled with pulse functions), tasks also consume a specified amount of reservoir resources at their start and produce a specified amount at their end (modeled with stepAtStart and stepAtEnd). The objective is to minimize the makespan. We experiment on 55 instances with 15 tasks adapted from (Koné et al. 2013). A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 4, MaxiCP solves more instances than Gecode and roughly the same set of instances as CPOptimizer. A backtrack analysis between CPOptimizer and MaxiCP reveals that the number of backtracks is nearly identical, suggesting that CPOptimizer is simply faster on this problem, as both solvers achieve similar levels of filtering on this problem.
SMIC
The Single Machine with Inventory Constraints (SMIC) problem was introduced in (Davari et al. 2020). Each task has a release date, a fixed duration, and a positive or negative inventory consumption that occurs at the start of the activity (modeled with stepAtStart). The inventory starts at a given value and must stay within a specified range. Activities cannot overlap in time. The objective is to minimize the makespan. We experiment on instances with 10 tasks from (Davari et al. 2020). The non-overlap constraint was enforced using the generalized cumulative constraint in MaxiCP (modeled using pulse) and Gecode. A time limit of 600s is set per instance for finding and proving optimality. As shown in Figure 4 CPOptimizer solves more instances. A backtrack analysis showed that CPOptimizer requires significantly fewer backtracks on this problem, while Gecode and MaxiCP are not on par for this problem in terms of time and number of backtracks. Unfortunately the results for CPOptimizer for this problem are biased as it appears to reformulate the cumulative function for modeling the non-overlap into a standard disjunctive constraint for which strong dedicated filtering exist such as (Vil´ım 2004).
MESP
The Maximum Energy Scheduling Problem (MESP) involves optional tasks with variable durations and resource demands, which can be positive or negative, on a single resource with fixed capacity. The objective is to maximize the total positive energy consumption of the resources while ensuring that the resource capacity is not exceeded at any time. All filtering rules are beneficial for this problem, making it the most relevant benchmark for evaluating the proposed filtering algorithm. We generated 60 instances with a number of tasks ranging between 6 and 12800. Given the large number of tasks in some instances, this experiment is designed to test the scalability of the filtering. A time limit of 2,000s and a memory limit of 16GB is set per instance for finding a feasible solution.
As can be observed in Figure 4, our approach performs very well with more instances solved. Both Gecode and CP Optimzer fail to solve largest instances due to timeouts and memory limits respectively. A backtrack analysis shows that our approach avoids backtracking, unlike the other solvers, and that CP Optimizer requires fewer backtracks than Gecode although it solves less instances.
Figure 4: Number of solved instances in function of time for the RCPSP-CPR, SMIC and MESP Problems.
Conclusion
We explained how the GeneralizedCumulative constraint can be used to implement modeling paradigm of cumulative functions for scheduling problems. We also introduced a time-tabling filtering algorithm for the GeneralizedCumulative constraint. The experimental results show that the proposed approach is scalable and competitive with other solvers on three cumulative scheduling problems with varying characteristics.

### v2

@@
 Implementing Cumulative Functions with Generalized Cumulative Constraints
-Abstract
+Abstract.
 Modeling scheduling problems with conditional time intervals and cumulative functions has become a common approach when using modern commercial constraint programming solvers.
@@
 In this work, we present an implementation of this modeling approach using a single, generic global constraint called the Generalized Cumulative.
-We also introduce a novel time-table filtering algorithm designed to handle tasks defined on conditional time-intervals.
-Experimental results demonstrate that this approach, combined with the new filtering algorithm, performs competitively with existing solvers enabling the modeling of producer and consumer scheduling problems and effectively scales to large problems.
-Introduction
-The success of Constraint Programming as a modeling and solving technology for hard scheduling problems is well established (Laborie et al.
-2018).
-The modeling has been further enhanced with the introduction of conditional time-interval variables (Laborie and Rogerie 2008), which represent the execution of tasks that may or may not be executed.
-Such variables allow to conveniently model problems with alternative resources or optional tasks such as in (Kinable, Wauters, and Berghe 2014; Kumar et al.
-2018; Cappart and Schaus 2017; Kizilay, Eliiyi, and Van Hentenryck 2018).
-Building on conditional time-intervals, the same authors introduced an algebraical model for cumulative resources called cumulative functions (Laborie et al.
-2009).
-It provides a convenient way to model producer-consumer problems with optional tasks such as in (Liu, Aleman, and Beck 2018; Gedik et al.
-2018; Cappart et al.
-2018; Thomas and Schaus 2024).
-CPOptimizer (Laborie et al.
-2018) and OptalCP (Vilím and Pons 2023) are the only two solvers that fully support conditional time-interval variables and cumulative functions and this modeling paradigm is not supported by existing open-source solvers [^1] API languages such as Minizinc also support optional variables at the modeling layer (Mears et al.
-2014).
-We propose an implementation of this modeling paradigm with lower and upper-bound restrictions on cumulative functions using a generalized cumulative constraint that supports negative heights (Beldiceanu and Carlsson 2002).
+We also introduce a novel time-table filtering algorithm specifically designed to handle tasks defined on conditional time-intervals.
+Experimental results demonstrate that this approach, combined with the new filtering algorithm, performs competitively with existing solvers enabling the modeling of producer and consumer scheduling problems and effectively scales to large-scale problems.
+1. Introduction
+The success of Constraint Programming as a modeling and solving technology for hard scheduling problems is well established (laborie2018ibm).
+The modeling flexibility has been further enhanced with the introduction of conditional time-interval variables (LaborieR08), which represent the execution of tasks that may or may not be executed.
+Such variables allow to conveniently model problems with alternative resources or optional tasks such as in (kinable2014concrete; kumar2018station; cappart2017rescheduling; kizilay2018constraint).
+Building on top of conditional time-intervals, the same authors introduced an algebraical model for cumulative resources called cumulative functions (LaborieRSV09).
+It provides an efficient and convenient way to model producer-consumer problems with optional tasks such as in (liu2018modelling; gedik2018constraint; cappart2018constraint; thomas2024constraint).
+To the best of our knowledge, CPOptimizer (laborie2018ibm) and OptalCP (OptalCP) are the only two (commercial) solvers that fully support conditional time-interval variables and cumulative functions and this modeling paradigm is not supported by existing open-source solvers [^1] API languages such as Minizinc also support optional variables at the modeling layer (optionminizinc).
+We propose an implementation of this modeling paradigm with lower and upper-bound restrictions on cumulative functions using a generalized cumulative constraint that supports negative heights (BeldiceanuC02).
 This more general form of the cumulative constraint allows variable and negative resource consumptions as well as optional tasks.
-We introduce a new filtering algorithm for this constraint that is based on time-tabling (Gay, Hartert, and Schaus 2015a; Fahimi, Ouellet, and Quimper 2018; Letort, Carlsson, and Beldiceanu 2015) and does additional filtering on the height and length of tasks.
+We introduce a new filtering algorithm for this constraint that is based on time-tabling (GayHS15; FahimiOQ18; LetortCB15) and does additional filtering on the height and length of tasks.
 The algorithm proceeds in two steps.
-First, it uses the Profile data structure recently introduced by (Gingras and Quimper 2016) and builds the optimistic/pessimistic profiles of resources produced/consumed by tasks.
+First, it uses the Profile data structure recently introduced by (GingrasQ16) and builds the optimistic/pessimistic profiles of resources produced/consumed by tasks.
 It then processes each task to prune its bounds by comparing it to the profile.
 The experimental results show that the newly introduced algorithm performs well compared to existing ones, particularly in large-scale scheduling problems solved with a greedy search.
-Modeling with Cumulative Functions
-Conditional Time-Intervals
-The domain of a conditional time-interval variable can be expressed as a subset of {}$\{\bot\}\cup\{[s,e)\mid s\leq e,s,e\in\mathbb{Z}\}$ (Laborie and Rogerie 2008).
+2. Modeling with Cumulative Functions
+2.1. Conditional Time-Intervals
+The domain of a conditional time-interval variable can be expressed as a subset of $\{\bot\}\cup\{[s,e)\mid s\leq e,s,e\in\mathbb{Z}\}$ (LaborieR08).
 It is fixed if $x=\bot$ (the interval is not executed) or $x=[s,e)$ (the interval starts at $s$ and ends at $e$).
@@
   • $p_{i}\in\{true,false\}$ is the execution status ($false=\bot$).
-  • $s_{i}\in[\underline{s}_{i},\overline{s}_{i}]$ and $e_{i}\in[\underline{e}_{i},\overline{e}_{i}]$ are the start/end variables.
+  • $s_{i}\in[\underline{s}_{i},\overline{s}_{i}]$ and $e_{i}\in[\underline{e}_{i},\overline{e}_{i}]$ are the (non negative) start/end variables.
   • $d_{i}\in[\underline{d}_{i},\overline{d}_{i}]$ is the (non negative) duration variable.
-A bound consistent filtering on the relation $s_{i}+d_{i}=e_{i}$ is maintained atomically on the corresponding attributes, and the domain of a time-interval variable becomes $\bot$ whenever one attribute range ($s,d$ or $e$) becomes empty which may trigger an inconsistency if $p_{i}=true$.
+A bound consistent filtering on the relation $s_{i}+d_{i}=e_{i}$ is maintained atomically on the corresponding attributes, and the domain of a time-interval variable becomes $\bot$ whenever one attribute range ($s,d$ or $e$) becomes empty which may trigger an inconsistency if $p_{i}=\mathit{true}$.
 The whole domain of a conditional time-interval $i\in T$ is thus encoded by a tuple of variables $x_{i}=\langle s_{i},d_{i},e_{i},p_{i}\rangle$.
-Cumulative Functions
-For modeling cumulative problems, (Laborie et al.
-2009) introduced the idea of cumulative function expressions.
+2.2. Cumulative Functions
+For modeling cumulative problems, (LaborieRSV09) introduced the idea of cumulative function expressions.
 They allow the expression of step-wise integer functions of conditional time-interval variables.
-The contribution of a single time-interval variable $i\in T$ is represented by an elementary cumulative function receiving an interval $x_{i}$ and a non-negative consumption value $c_{i}$ or a non-negative height range $[\underline{c}_{i},\overline{c}_{i}]$.
-The two main [^2] Variants exist such as stepAtEnd.
-elementary functions are pulse and stepAtStart (Laborie et al.
-2009) as represented in Figure 1.
-Figure 1: Elementary cumulative functions.
+The contribution of a single time-interval variable $i\in T$ is represented by an elementary cumulative function receiving an interval $x_{i}$ and a non-negative consumption value $c_{i}$ or a non-negative height range $\underline{c}_{i},\overline{c}_{i}]$.
+The elementary functions are pulse, stepAtStart and stepAtEnd (LaborieRSV09) as represented in Figure 1.
+Figure 1. Elementary cumulative functions.
 Cumulative functions (elementary or not) can then be combined with the plus/minus operators to form a more complex function representing the addition or subtraction of two functions.
@@
 Example 0.
-Three tasks $A$, $B$ and $C$ contribute to a cumulative function $f=stepAtStart(A,2)\allowbreak-\allowbreak(pulse(B,1)\allowbreak+\allowbreak stepAtEnd(C,1))$.
+Three tasks $A$, $B$ and $C$ contribute to a cumulative function $f=\mathit{stepAtStart}(A,2)\allowbreak-\allowbreak(pulse(B,1)\allowbreak+\allowbreak\mathit{stepAtEnd}(C,1))$.
 The resulting AST and cumulative function are represented in Figure 2.
@@
 A dotted line represents the final cumulative function $f$.
-Figure 2: Example 1: AST (left) and cumul. function (right).
-A cumulative function can be flattened using Algorithm 1 to extract a set of cumulative tasks (tasks with their positive or negative heights).
+Figure 2. Example 1: AST (left) and cumul. function (right).
+A cumulative function can be flattened using Algorithm 1 to extract a set of cumulative tasks (tasks with their positive or negative heights represented by a tuple $(x_{i},c_{i})$).
 This algorithm traverses the AST and collects the activities at the leaf nodes, reversing the height signs for the activities flattened on the right branch of a minus node.
-Note that whenever encountering a $stepAtStart(x,c)$ leaf node, the cumulative task collected is defined on a fresh task interval $x^{\prime}$ related to the original one with: $x^{\prime}.s=x.s$ (same start as $x$), $x^{\prime}.p=x.p$ (same execution status), and $x^{\prime}.e=Horizon$ (ending at the time horizon) to reflect the definition of a step function [^3] Other cases such as $stepAtEnd(x,c)$ are defined similarly..
-1 if $n=\text{Pulse}(x,c)$ then  $\{(x,c)\}$
-2 else if $n=\text{StepAtStart}(x,c)$ then  $\{(x^{\prime},c)\}$
-3 else if $n=\text{Plus}(n_{l},n_{r})$ then
-4 $\text{flatten}(n_{l})\cup\text{flatten}(n_{r})$
-5else if $n=\text{Minus}(n_{l},n_{r})$ then
-6 $\text{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\text{flatten}(n_{r})\}$
-7else  $error$ // Unknown pattern
-[…]
+Note that whenever encountering a $\mathit{stepAtStart}(x,c)$ leaf node, the cumulative task collected is defined on a fresh task interval $x^{\prime}$ related to the original one with: $x^{\prime}.s=x.s$ (same start as $x$), $x^{\prime}.p=x.p$ (same execution status), and $x^{\prime}.e=Horizon$ (ending at the time horizon) to reflect the definition of a step function.
+Similarly, in the case of $\mathit{stepAtEnd}(x,c)$, a new task interval $x^{\prime}$ with $x^{\prime}.s=x.e$ (same start as end of $x$) is collected.
+Input: $n$: A node of the AST
+Output: Set of cumulative tasks
+1 if $n=\mathit{pulse}(x,c)$ then $\{(x,c)\}$
+2 else if $n=\mathit{stepAtStart}(x,c)$ then $\{(x^{\prime}\leftarrow\langle x.s,Horizon-x.s,Horizon,x.p\rangle,c)\}$
+3 else if $n=\mathit{stepAtEnd}(x,c)$ then $\{(x^{\prime}\leftarrow\langle x.e,Horizon-x.e,Horizon,x.p\rangle,c)\}$
+4 else if $n=\mathit{plus}(n_{l},n_{r})$ then $\mathit{flatten}(n_{l})\cup\mathit{flatten}(n_{r})$
+5 else if $n=\mathit{minus}(n_{l},n_{r})$ then $\mathit{flatten}(n_{l})\cup\{(x,-c)\mid(x,c)\in\mathit{flatten}(n_{r})\}$
+6 else errorerror // Unknown pattern
+Algorithm 1 flatten($n$)
 Example 0.
-Continuing example 1, $f$) collects the cumulative tasks $\{(A^{\prime},2),(B,-1),(C^{\prime},-1)\}$ illustrated on the right of Figure 2.
-Constraints on Cumulative Functions
+Continuing example 1, flatten($f$) collects the cumulative tasks $\{(A^{\prime},2),(B,-1),(C^{\prime},-1)\}$ illustrated on the right of Figure 2.
+2.3. Constraints on Cumulative Functions
 The height of a cumulative function can be restricted with a lower and upper bound at every time overlapping at least one executing time-interval: $\underline{C}\leq f\leq\overline{C}$.
-This constraint is called $alwaysIn(f,\underline{C},\overline{C})$ in (Laborie et al.
-2009).
+This constraint is called $alwaysIn(f,\underline{C},\overline{C})$ in (LaborieRSV09).
 It can be enforced by posting a GeneralizedCumulative constraint requiring a resource consumption between $\underline{C}$ and $\overline{C}$ for the set of cumulative tasks $T$ collected with the Flatten function.
@@
 Definition 0.
-The set $O=\{i\in T\mid p_{i}=\{true,false\}\}$ denotes the set of optional tasks, those for which it is not decided yet if they will execute or not on the resource.
-The set $R=\{i\in T\mid p_{i}=true\}$ denotes the set of required tasks, those which will surely be executed on the resource.
-The set $E=\{i\in T\mid p_{i}=false\}$ denotes the set of excluded tasks, those which will surely not be executed on the resource.
+The set $O=\{i\in T\mid p_{i}=\{\mathit{true},\mathit{false}\}\}$ denotes the set of optional tasks, those for which it is not decided yet if they will execute or not on the resource.
+The set $R=\{i\in T\mid p_{i}=\mathit{true}\}$ denotes the set of required tasks, those which will surely be executed on the resource.
+The set $E=\{i\in T\mid p_{i}=\mathit{false}\}$ denotes the set of excluded tasks, those which will surely not be executed on the resource.
 The sets $O$, $R$ and $E$ are disjoint and form a partition of the set $T$, i.e., $T=O\cup R\cup E$.
 Definition 0.
-[…]
-$\sum\limits_{i\in R\mid s_{i}\leq\tau<e_{i}}c_{i}\in[\underline{C},\overline{C}]\qquad\forall\tau\in\bigcup_{i\in R}[s_{i},e_{i})$ (1)
+The $\texttt{GeneralizedCumulative}(T,\underline{C},\overline{C})$ constraint enforces the cumulated heights of the tasks of $T$ to be within the fixed capacity range $\underline{C},\overline{C}]$ for every time at which at least one task of $T$ is executed:
+(1) $\sum\limits_{i\in R\mid s_{i}\leq\tau<e_{i}}c_{i}\in[\underline{C},\overline{C}]\qquad\forall\tau\in\bigcup_{i\in R}[s_{i},e_{i})$
 Restricted Time Interval Scope
@@
 Consistency
-Enforcing global bound-consistency for the cumulative constraint is NP-Hard (Garey and Johnson 1979).
+Enforcing global bound-consistency for the cumulative constraint is NP-Hard (GareyJ79).
 However, enforcing it on the decomposition into sum constraints at every point in the horizon, as in (1), is not NP-Hard.
@@
 This filtering is generally referred to as time-tabling filtering.
-Intuitively, it works by aggregating all the fixed parts (when $\underline{s}_{i}<\overline{e}_{i}$) to postpone task execution.
-Timetabling Algorithm
-The Profile Data Structure
-The Profile data structure is an aggregation of adjoined rectangles of different lengths and heights introduced in (Gingras and Quimper 2016; Gay, Hartert, and Schaus 2015a) to record the resource utilization of tasks.
+Intuitively, it works by aggregating all the fixed parts (when $\underline{s}_{i}<\overline{e}_{i}$) to create a consumption profile.
+It then postpones the earliest execution of a task whenever it detects that starting it at this time would overload the resource capacity.
+3. Timetabling Algorithm
+This section first introduces the Profile data structure then explains the time-tabling filtering algorithm for filtering the task interval attributes.
+3.1. The Profile Data Structure
+The Profile data structure is an aggregation of adjoined rectangles of different lengths and heights introduced in (GingrasQ16; GayHS15) to record the resource utilization of tasks.
 The end of a rectangle is the beginning of the next one.
@@
 ending) time of the rectangle and other components that contain information relative to the resource consumption for the duration of the rectangle.
-These time points are sorted in increasing order of time and are kept in a linked list structure called Profile.
-An original idea of (Gingras and Quimper 2016) is to have pointers $tp(\underline{s}_{i})$, $tp(\underline{e}_{i})$ and $tp(\overline{e}_{i})$ linking to the time point associated with $\underline{s}_{i}$, $\underline{e}_{i}$, and $\overline{e}_{i}$ respectively.
-Timetable Filtering
-[…]
+These time points are sorted in increasing order of time and are kept in a linked list structure called the Profile.
+An original idea of (GingrasQ16) is to have pointers $tp(\underline{s}_{i})$, $tp(\underline{e}_{i})$ and $tp(\overline{e}_{i})$ linking to the time point associated with $\underline{s}_{i}$, $\underline{e}_{i}$, and $\overline{e}_{i}$ respectively.
+3.2. Timetable Filtering
+To check and adjust the attributes of tasks, the Profile data structure from (GingrasQ16) is extended to compute both a minimum and maximum profile range at each time point.
+In our adaptation of the Profile data structure (hereafter referred to as timeline and noted $P$), a time point $tp\in P$ consists of a tuple $\langle\mathit{time},\underline{P},\overline{P},\mathit{\#fp}\rangle$, where $\mathit{time}$ corresponds to the start time of the time point, $\underline{P}$ (resp.
+$\overline{P}$) to the minimum (resp.
+maximum) resource profile of tasks over time and $\mathit{\#fp}$ to the number of fixed parts of tasks that overlap over $\mathit{time}$.
 Intuitively, the minimum profile is obtained by considering the upper bound of the negative contribution and the lower bound of the positive contribution of each task.
@@
 maximum $\overline{P}$) resource profile at time $t$ is defined as
-$\underline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\min(\underline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\max(\underline{c}_{i},0)$ (2)
-$\overline{P}_{t}=\sum\limits_{\mathclap{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}}\max(\overline{c}_{i},0)+\sum\limits_{\mathclap{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}}\min(\overline{c}_{i},0)$ (3)
+(2) $\underline{P}_{t}=\sum\limits_{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}\min(\underline{c}_{i},0)+\sum\limits_{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}\max(\underline{c}_{i},0)$
+(3) $\overline{P}_{t}=\sum\limits_{i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}}\max(\overline{c}_{i},0)+\sum\limits_{i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}}\min(\overline{c}_{i},0)$
 where $i\in O\cup R\mid\underline{s}_{i}\leq t<\overline{e}_{i}$ corresponds to the set of all optional and required tasks whose time window overlaps the time $t$ and $i\in R\mid\overline{s}_{i}\leq t<\underline{e}_{i}$ corresponds to the set of all required tasks that have a fixed part that overlaps the time $t$.
@@
 Second, each task is checked during its entire possible span in terms of profile and capacity to adjust its domain.
-Timeline Initialization and Consistency Check
+3.3. Timeline Initialization and Consistency Check
 The function initializeTimeline (Algorithm 2) initializes the timeline, links the corresponding time points to tasks, and computes the minimum and maximum profile of the resource utilization.
-It receives as input the set of tasks $T$.
+It receives as input the set of tasks $T$ and the bounds of the resource capacity $\underline{C},\overline{C}]$.
 First, events are generated by iterating over the tasks that are not absent (Line 2).
@@
 For tasks that are present and with a fixed part, two additional events are generated for the maximum start time ($\overline{s}_{i}$) and the minimum end time ($\underline{e}_{i}$) of the task.
-1 ${}events\leftarrow\{\}$
-2 for $i\in T\mid i\in O\cup R$ do
-3 $events\leftarrow events\cup\{\underline{s}_{i},\overline{e}_{i}\}$
-4 if $i\in R\wedge\overline{s}_{i}<\underline{e}_{i}$ then
-5    $events\leftarrow events\cup\{\overline{s}_{i},\underline{e}_{i}\}$
-[…]
-7 $tp\leftarrow new\,tp(\min(events.time))$
-8 $P\leftarrow\{tp\}$
-9 for $e\in events$ sorted by time do
-10 $i\leftarrow e.task$
-11 if $tp.time<e.time$ then
-12    if #$\#fp>0\wedge(\underline{P}>\overline{C}\vee\overline{P}<\underline{C})$ then
-13       return Fail
-14
-15    $tp\leftarrow new\,tp(e.time)$
-16    $P.append(tp)$
-17
-18 if $e=\underline{s}_{i}$ then
-19    $\underline{P}\leftarrow\underline{P}+\min(\underline{c}_{i},0)$
-20    $\overline{P}\leftarrow\overline{P}+\max(\overline{c}_{i},0)$
-21    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
-22    $tp(\underline{s}_{i})\leftarrow tp$
-23
-24 if $e=\overline{e}_{i}$ then
-25    $\underline{P}\leftarrow\underline{P}-\min(\underline{c}_{i},0)$
-26    $\overline{P}\leftarrow\overline{P}-\max(\overline{c}_{i},0)$
-27    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
-28    $tp(\overline{e}_{i})\leftarrow tp$
-29
-30 if $e=\overline{s}_{i}$ then
-31    $\underline{P}\leftarrow\underline{P}+\max(\underline{c}_{i},0)$
-32    $\overline{P}\leftarrow\overline{P}+\min(\overline{c}_{i},0)$
-33    #$\#fp\leftarrow\#fp+1$
-34    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
-35
-36 if $e=\underline{e}_{i}$ then
-37    $\underline{P}\leftarrow\underline{P}-\max(\underline{c}_{i},0)$
-38    $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
-39    #$\#fp\leftarrow\#fp-1$
-40    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\#fp\leftarrow\#fp$
-41
-return $P$
-[…]
 Then, these events are processed by order of increasing time to initialize the timeline $P$ (Line 2).
@@
 When processing a start min ($\underline{s}$) or end max ($\overline{e}$) event, the current time point is linked to the event at Lines 2 and 2 to retrieve the time points corresponding to the start and end of tasks in constant time later.
-[…]
+Three accumulators are used to keep track of the minimum ($\underline{P}$) and maximum ($\overline{P}$) profile as well as the number of fixed parts of tasks overlapping the current time point ($\mathit{\#fp}$).
+Depending on the nature of the event, the positive or negative contribution of its associated task is either added to (in case of $\underline{s}$ or $\overline{s}$) or subtracted (in case of $\overline{e}$ or $\underline{e}$) from the accumulators and the current time point $tp$ is updated.
 As the timeline is initialized, the consistency of the minimum and maximum profiles is checked for each time point at Line 2.
-Once a time point is completed (when creating the next time point), we ensure that if at least one task is guaranteed to execute during the time point (#$\#fp>0$), the minimum profile $\underline{P}$ is not above the maximum capacity of the resource $\overline{C}$ and the maximum profile $\overline{P}$ is not below the minimum capacity of the resource $\underline{C}$.
+Once a time point is completed (when creating the next time point), we ensure that if at least one task is guaranteed to execute during the time point ($\mathit{\#fp}>0$), the minimum profile $\underline{P}$ is not above the maximum capacity of the resource $\overline{C}$ and the maximum profile $\overline{P}$ is not below the minimum capacity of the resource $\underline{C}$.
 The worst-case complexity in time of the function initializeTimeline is $\mathcal{O}(n\log(n))$ and is due to the sorting of the events used in the loop of Line 2.
 Example 1 illustrates the profile initialization for three tasks.
+$\mathit{events}\leftarrow\{\}$
+// Initialize the set of events
+1 for $i\in T\mid i\in O\cup R$ do
+ $\mathit{events}\leftarrow\mathit{events}\cup\{\underline{s}_{i},\overline{e}_{i}\}$
+ // Add events for non-absent tasks
+2 if $i\in R\wedge\overline{s}_{i}<\underline{e}_{i}$ then
+    $\mathit{events}\leftarrow\mathit{events}\cup\{\overline{s}_{i},\underline{e}_{i}\}$
+    // Add events for required tasks with fixed part
+3
+$\underline{P}\leftarrow 0$; $\overline{P}\leftarrow 0$
+// Lower/Upper bound of the profile at current time point
+$\mathit{\#fp}\leftarrow 0$
+// Number of overlapping fixed parts of current time point
+4
+$tp\leftarrow\textbf{new}\,tp(\min(\mathit{events}.\mathit{time}))$
+// Current time point
+$P\leftarrow\{tp\}$
+// All time points created so-far
+5
+6for $e\in\mathit{events}$ sorted by time do
+7 $i\leftarrow e.task$
+8 if $tp.\mathit{time}<e.\mathit{time}$ then
+9    if $\mathit{\#fp}>0\wedge(\underline{P}>\overline{C}\vee\overline{P}<\underline{C})$ then
+       return Fail
+       // Capacity violation detected
+10
+    $tp\leftarrow\textbf{new}\,tp(e.\mathit{time})$
+    // New time point
+11    $P.append(tp)$
+12
+13 if $e=\underline{s}_{i}$ then // Process earliest start event
+    $\underline{P}\leftarrow\underline{P}+\min(\underline{c}_{i},0)$
+    // Add pessimistic negative contribution
+    $\overline{P}\leftarrow\overline{P}+\max(\overline{c}_{i},0)$
+    // Add optimistic positive contribution
+    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$; $tp(\underline{s}_{i})\leftarrow tp$
+    // Update tp and link task to tp
+14
+15 if $e=\overline{e}_{i}$ then // Process latest end event
+    $\underline{P}\leftarrow\underline{P}-\min(\underline{c}_{i},0)$
+    // Remove pessimistic negative contribution
+    $\overline{P}\leftarrow\overline{P}-\max(\overline{c}_{i},0)$
+    // Remove optimistic positive contribution
+    $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$; $tp(\overline{e}_{i})\leftarrow tp$
+    // Update tp and link task to tp
+16
+17 if $p_{i}=\mathit{true}\wedge\overline{s}_{i}<\underline{e}_{i}$ then
+18    if $e=\overline{s}_{i}$ then // Process start of fixed parts event
+       $\underline{P}\leftarrow\underline{P}+\max(\underline{c}_{i},0)$
+       // Add pessimistic positive contribution
+       $\overline{P}\leftarrow\overline{P}+\min(\overline{c}_{i},0)$
+       // Add optimistic negative contribution
+       $\mathit{\#fp}\leftarrow\mathit{\#fp}+1$
+       // Increment overlapping fixed part counter
+19       $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$
+20
+21    if $e=\underline{e}_{i}$ then // Process end of fixed parts event
+       $\underline{P}\leftarrow\underline{P}-\max(\underline{c}_{i},0)$
+       // Remove pessimistic positive contribution
+       $\overline{P}\leftarrow\overline{P}-\min(\overline{c}_{i},0)$
+       // Remove optimistic negative contribution
+       $\mathit{\#fp}\leftarrow\mathit{\#fp}-1$
+       // Decrement overlapping fixed part counter
+22       $tp.\underline{P}\leftarrow\underline{P}$; $tp.\overline{P}\leftarrow\overline{P}$; $tp.\mathit{\#fp}\leftarrow\mathit{\#fp}$
+23
+return $P$
+// Return the complete timeline
+Algorithm 2 initializeTimeline($T,\underline{C},\overline{C}$)
 Example 0.
 Let us consider three tasks :
-$A$
-: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=[1,2],\allowbreak p=true\rangle$;
-$B$
-: $\langle s=[2,4],\allowbreak d=[3,4],\allowbreak e=[5,7],\allowbreak c=2,\allowbreak p=true\rangle$;
-$C$
-: $\langle s=[3,8],\allowbreak d=[1,3],\allowbreak e=[4,9],\allowbreak c=[-2,1],\allowbreak p=\{true,false\}\rangle$;
-and a resource $C$ of capacity bounds [0,1][0,1][ 0 , 1 ].
-The time points and their attributes are shown in Table 1.
-Figure 3 shows a representation of the tasks and the resulting profile ranges.
-Note that the events $\overline{s}_{c}$ and $\underline{e}_{c}$ are not considered as task $C$ has no fixed part.
-[…]
+AA:
+: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=[1,2],\allowbreak p=\mathit{true}\rangle$;
+$B$:
+: $\langle s=[2,4],\allowbreak d=[3,4],\allowbreak e=[5,7],\allowbreak c=2,\allowbreak p=\mathit{true}\rangle$;
+$C$:
+: $\langle s=[3,8],\allowbreak d=[1,3],\allowbreak e=[4,9],\allowbreak c=[-2,1],\allowbreak p=\{\mathit{true},\mathit{false}\}\rangle$;
+and a resource $C$ of capacity bounds $[0,1]$. The time points and their attributes are shown in Table 1.
+time event(s) $\underline{P}$ $\overline{P}$ $\mathit{\#fp}$
 0 $\underline{s}_{a}$ 0 2 0
@@
 2 $\underline{s}_{b}$ 1 4 1
-3 $\underline{e}_{a},\underline{s}_{c}$2 5 0
+3 $\underline{e}_{a},\underline{s}_{c}$ -2 5 0
 4 $\overline{e}_{a},\overline{s}_{b}$ 0 3 1
-5 $\underline{e}_{b}$2 3 0
-7 $\overline{e}_{b}$2 1 0
+5 $\underline{e}_{b}$ -2 3 0
+7 $\overline{e}_{b}$ -2 1 0
 9 $\overline{e}_{c}$ 0 0 0
-Table 1: Time points computed from the tasks of Example 3.
-Figure 3: Representation of the tasks of Example 3 with the resulting minimum ($\underline{P}$) and maximum ($\overline{P}$) profile.
-Filtering
+Table 1. Time points computed from the tasks of Example 1.
+Figure 3 shows a representation of the tasks and the resulting profile ranges. Note that the events $\overline{s}_{c}$ and $\underline{e}_{c}$ are not considered as task $C$ has no fixed part.
+Figure 3. Representation of the tasks of Example 1 with the resulting minimum ($\underline{P}$) and maximum ($\overline{P}$) profile.
+3.4. Filtering
 Once the timeline is initialized and its consistency verified, the profile range (minimum and maximum profile of tasks at any time point) is used to filter the time windows, capacity, and duration of tasks.
@@
 Formally, the rule is defined as:
-$\forall tp\in P\,|\,\underline{s}_{i}\leq tp.time<\min(\overline{s}_{i},\underline{e}_{i}),$
-$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (4)
-$\Rightarrow\underline{s}_{i}\geq tp.next.time$
-$\forall tp\in P\,|\,\max(\overline{s}_{i},\underline{e}_{i})\leq tp.time<\overline{e}_{i},$
-$tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ (5)
-$\Rightarrow\overline{e}_{i}\leq tp.prev.time$
+(4) $\forall tp\in P\,|\,\underline{s}_{i}\leq tp.\mathit{time}<\min(\overline{s}_{i},\underline{e}_{i}),\quad tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}\Rightarrow\underline{s}_{i}\geq tp.\mathit{next}.\mathit{time}$
+(5) $\forall\mathit{tp}\in P\,|\,\max(\overline{s}_{i},\underline{e}_{i})\leq\mathit{tp.\mathit{time}}<\overline{e}_{i},\quad tp.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}\Rightarrow\overline{e}_{i}\leq\mathit{tp.\mathit{prev}.\mathit{time}}$
 Note that as tasks are modeled with interval variables, the rule $e_{i}=s_{i}+d_{i}$ is enforced internally in the variable.
@@
 The second rule detects if the execution of a task $i\in T$ at a time point $tp$ is necessary to avoid a violation of the resource capacity range at this time point.
-If this is the case, the task is adjusted to make sure it is present during the time point: 1) if the task is optional it is set to present; 2) its time window is adjusted such that it executes over the whole time point; 3) its height is adjusted to avoid the profile range violation.
+If this is the case, the task is adjusted to make sure it is present during the time point:
+  1. (1)
+    if the task is optional, it is set to present;
+  2. (2)
+    its time window is adjusted such that it executes over the whole time point;
+  3. (3)
+    its height is adjusted to avoid the profile range violation.
 In formal terms, the rule is defined as:
-$\forall tp\in P\,|\,tp.time\in[\underline{s}_{i},\overline{e}_{i})\wedge tp.\#fp>0,$
-$tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C}$
-$\Rightarrow\begin{cases}&p_{i}\leftarrow true\\ &\overline{s}_{i}\leq tp.time\\ &\underline{e}_{i}\geq tp.next.time\\ &\underline{c}_{i}\geq deficit\text{ if }deficit>0\\ &\overline{c}_{i}\leq overload\text{ if }overload<0\end{cases}$ (6)
-where $deficit=\underline{C}-(\overline{P}-\max(\overline{c}_{i},0))$ and $overload=\overline{C}-(\underline{P}-\min(\underline{c}_{i},0))$ are the deficit and overload of height needed for the profile range to intersect the capacity range.
+$\forall\mathit{tp}\in P\,|\,\mathit{tp.\mathit{time}}\in[\underline{s}_{i},\overline{e}_{i})\wedge\mathit{tp.\mathit{\#fp}}>0,$
+(6) $\mathit{tp}.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\mathit{tp}.\overline{P}-\max(\overline{c}_{i},0)<\underline{C}\Rightarrow\begin{cases}&p_{i}\leftarrow\mathit{true}\\ &\overline{s}_{i}\leq\mathit{tp.\mathit{time}}\\ &\underline{e}_{i}\geq\mathit{tp.\mathit{next}.\mathit{time}}\\ &\underline{c}_{i}\geq\mathit{deficit}\text{ if }\mathit{deficit}>0\\ &\overline{c}_{i}\leq\mathit{overload}\text{ if }\mathit{overload}<0\end{cases}$
+where $\mathit{deficit}=\underline{C}-(\overline{P}-\max(\overline{c}_{i},0))$ and $\mathit{overload}=\overline{C}-(\underline{P}-\min(\underline{c}_{i},0))$ are the deficit and overload of height needed for the profile range to intersect the capacity range.
 Height
@@
 Two cases are possible:
-1). If the task has a fixed part, its height is checked and adjusted directly at each time point of the fixed part:
-$\forall tp\in P\mid\overline{s}_{i}\leq tp.time<\underline{e}_{i},$
-$\overline{c}_{i}\leq\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$ (7)
-$\underline{c}_{i}\geq\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
+  1. (1)
+    If the task has a fixed part, its height is checked and adjusted directly at each time point of the fixed part:
+    (7) $\forall tp\in P\mid\overline{s}_{i}\leq tp.\mathit{time}<\underline{e}_{i},\quad\overline{c}_{i}\leq\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge\underline{c}_{i}\geq\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
     Note that if a task is present ($i\in R$), its minimum positive ($\max(\underline{c}_{i},0)$) and negative ($\min(\overline{c}_{i},0)$) contributions have been used to compute the profile range and thus these values must be subtracted in the above equation.
-2). If the task has no fixed part, the maximum available height over the minimum overlapping interval of the task is used to adjust its height:
-$\overline{c}_{i}\leq\max_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))\,\wedge$
-$\underline{c}_{i}\geq\min_{\mathclap{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.time\leq\overline{s}_{i}}}\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$ (8)
-The minimum overlapping interval was introduced by (Gay, Hartert, and Schaus 2015b) and is defined for a task $i\in T$ as the interval $[\underline{e}_{i}-1,\overline{s}_{i}]$.
+  2. (2)
+    If the task has no fixed part, the maximum available height over the minimum overlapping interval of the task is used to adjust its height:
+    (8) $\overline{c}_{i}\leq\max_{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.\mathit{time}\leq\overline{s}_{i}}\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
+    (9) $\underline{c}_{i}\geq\min_{tp\in P\mid\underline{e}_{i}-1<tp.end\wedge tp.\mathit{time}\leq\overline{s}_{i}}\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
+    The minimum overlapping interval was introduced by (gay2015time) and is defined for a task $i\in T$ as the interval $\underline{e}_{i}-1,\overline{s}_{i}]$.
 Intuitively, it corresponds to the smallest interval such that, no matter its start time, length or end time, the task executes during at least one time unit of this interval.
 Length
-The last rule is inspired by (Ouellet and Quimper 2019).
+The last rule is inspired by (ouelletProcessing).
 It adjusts the maximum duration $\overline{d}$ of the tasks that do not have a fixed part (for the tasks with a fixed part, the adjustment would be redundant with the Forbid rule).
 The principle is to find the longest time span during which the task can be scheduled without a resource capacity violation.
-The function timetabling of Algorithm 3 checks these rules for each task $i\in T$.
-This algorithm receives as input the set of tasks $T$ and the bounds of the resource capacity $[\underline{C},\overline{C}]$.
+Again, we rely on the internal propagation rules of the interval variables to set the task as absent if the adjustment would empty its domain and to update the $\underline{s}_{i}$ and $\overline{e}_{i}$ attributes following the relations $\underline{s}_{i}\geq\underline{e}_{i}-\overline{d}_{i}$ and $\overline{e}_{i}\leq\overline{s}_{i}+\overline{d}_{i}$.
+Formally, the rule is written as:
+$\overline{d}_{i}\leq\max_{[a,b)\in A}b-a$
+(10) where
+$A=\left\{[a,b)\subseteq[\underline{s}_{i},\overline{e}_{i}]\mid\forall tp\in P\text{ with }tp.\mathit{time}\in[a,b),\;tp.\underline{P}+\max(\underline{c}_{i},0)\leq\overline{C}\wedge tp.\overline{P}+\min(\overline{c}_{i},0)\geq\underline{C}\right\}$
+Intuitively, $A$ is the set of intervals where the task can be scheduled without underloading the minimum and overloading maximum capacity.
+Algorithm
+The function timetabling of Algorithm 3 enforces these rules for each task $i\in T$.
+This algorithm receives as input the set of tasks $T$ and the bounds of the resource capacity $\underline{C},\overline{C}]$.
 The algorithm iterates over all non-fixed tasks at Line 3.
-1 $P\leftarrow\text{initializeTimeLine}(T)$
-2 $O\leftarrow\{i\in T\mid p_{i}=\{true,false\}\}$
-3 $R\leftarrow\{i\in T\mid p_{i}=true\}$
+Each task is processed in three steps:
+First (step 1), the Forbid rule is used to adjust the minimum start time of the task by iterating forward over the time points in the interval $\underline{s}_{i},\min(\overline{s}_{i},\underline{e}_{i}))$ in the loop at Line 3.
+Second (step 2), the Forbid rule is used to adjust the maximum end time of the task by iterating backward over the interval $\max(\overline{s}_{i},\underline{e}_{i}),\overline{e}_{i})$ in the loop at Line 3.
+The last step iterates over the time points of the remaining part of the time window.
+It depends if the task has a fixed part or not.
+In the former case (step 3), the loop at Line 3 iterates over the time points in the fixed part of the task and adjusts its height following the 1st case of the Height rule.
+In the case where the task has no fixed part (step 3’), the loop at Line 3 is executed.
+The height of the task is adjusted according to the 2nd case of the Height rule and its maximum length is adjusted following the Length rule.
+These three steps are illustrated in Figure 4.
+Figure 4. Timetabling processing order for a task with a fixed part (top) and one without (bottom). Those steps (1, 2, 3, and 3’) are also referenced in Algorithm 3 comments.
+$P\leftarrow\text{initializeTimeLine}(T)$
+// Compute the profile
+$O\leftarrow\{i\in T\mid p_{i}=\{\mathit{true},\mathit{false}\}\}$; $R\leftarrow\{i\in T\mid p_{i}=\mathit{true}\}$
+// Optional/Required tasks
 1 for $i\in O\cup R$ do
 2 if $i$ is fixed then continue
-6 $tpf\leftarrow tp(\underline{s}_{i})$
-7 while $tpf.time<\min(\overline{s}_{i},\underline{e}_{i})$ do
-8    if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
-9       $\underline{s}_{i}\leftarrow\max(tpf.next.time,\underline{s}_{i})$
+ $\mathit{tpf}\leftarrow tp(\underline{s}_{i})$
+ // Get the time point at minimum start
+3 while $\mathit{tpf}.\mathit{time}<\min(\overline{s}_{i},\underline{e}_{i})$ do // Step 1 (forward): adjust minimum start time
+    // Check if task can start at this time point, if not push its start
+4    if $\mathit{tpf}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpf}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $\underline{s}_{i}\leftarrow\max(\mathit{tpf}.\mathit{next}.\mathit{time},\underline{s}_{i})$
+5    else $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
+    $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
+    // Move to next time point
+6
+ $\mathit{tpb}\leftarrow tp(\overline{e}_{i}).\mathit{prev}$
+ // Get the time point just before maximum end
+7 while $\mathit{tpb}.\mathit{next}.\mathit{time}>\max(\overline{s}_{i},\underline{e}_{i})$ do // Step 2 (backward): adjust maximum end time
+    // Check if task can end after this time point, if not pull its end
+8    if $\mathit{tpb}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpb}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $\overline{e}_{i}\leftarrow\min(\mathit{tpb}.\mathit{time},\overline{e}_{i})$
+9    else $\text{checkIfMandatory}(i,\mathit{tpb},\underline{C},\overline{C})$
+    $\mathit{tpb}\leftarrow\mathit{tpb}.\mathit{prev}$
+    // Move to previous time point
 10
-11    else
-[…]
-13
-14    $tpf\leftarrow tpf.next$
+11 if $\overline{s}_{i}<\underline{e}_{i}$ then // Step 3: Adjust height (task with a fixed part)
+12    while $\mathit{tpf}.\mathit{time}<\underline{e}_{i}$ do // Scan over fixed part $\overline{s}_{i},\underline{e}_{i}-1,]$
+13       $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
+       $\underline{c}^{\prime}\leftarrow\underline{C}-(\mathit{tpf}.\overline{P}-\max(\overline{c}_{i},0))$; $\overline{c}^{\prime}\leftarrow\overline{C}-(\mathit{tpf}.\underline{P}-\min(\underline{c}_{i},0))$ // Available height range
+14       if $i\in R$ then
+          $\underline{c}^{\prime}\leftarrow\underline{c}^{\prime}+\min(\overline{c}_{i},0)$; $\overline{c}^{\prime}\leftarrow\overline{c}^{\prime}+\max(\underline{c}_{i},0)$ // Adjust for task’s own contribution P
 15
-16 $tpb\leftarrow tp(\overline{e}_{i}).prev$
-17 while $tpb.next.time>\max(\overline{s}_{i},\underline{e}_{i})$ do
-18    if $tpb.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpb.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
-19       $\overline{e}_{i}\leftarrow\min(tpb.time,\overline{e}_{i})$
-20
-21    else
-[…]
-23
-24    $tpb\leftarrow tpb.prev$
+       $\underline{c}_{i}\leftarrow\max(\underline{c}^{\prime},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{\prime},\overline{c}_{i})$
+       // Restrict min/max task consumption
+       $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
+       // Move to next time point
+16
+17 else // Step 3’: Adjust height and length (task without a fixed part)
+    $\overline{d}^{*}\leftarrow 0$
+    // Track maximum feasible duration
+    $s^{\prime}\leftarrow\underline{s}_{i}$
+    // Track start of current feasible interval
+    $\underline{c}^{*}\leftarrow\underline{C}-\mathit{tpf}.\mathit{prev}.\overline{P}+\max(\overline{c}_{i},0)$
+    // Min height over minimum overlapping interval
+    $\overline{c}^{*}\leftarrow\overline{C}-\mathit{tpf}.\mathit{prev}.\underline{P}+\min(\underline{c}_{i},0)$
+    // Max height over minimum overlapping interval
+18    while $\mathit{tpf}.\mathit{time}<\overline{s}_{i}$ do // Scan over minimum overlapping interval $\underline{e}_{i}-1,\overline{s}_{i}]$
+19       $\overline{d}^{*}\leftarrow\max(\mathit{tpf}.\mathit{time}-s^{\prime},\overline{d}^{*})$
+20       if $\mathit{tpf}.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak\mathit{tpf}.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then $s^{\prime}\leftarrow\mathit{tpf}.\mathit{next}.\mathit{time}$
+21       // Reset startelse $\text{checkIfMandatory}(i,\mathit{tpf},\underline{C},\overline{C})$
+22       $\underline{c}^{*}\leftarrow\min(\underline{C}-\mathit{tpf}.\overline{P}+\max(\overline{c}_{i},0),\underline{c}^{*})$
+23       $\overline{c}^{*}\leftarrow\max(\overline{C}-\mathit{tpf}.\underline{P}+\min(\underline{c}_{i},0),\overline{c}^{*})$
+24       $\mathit{tpf}\leftarrow\mathit{tpf}.\mathit{next}$
 25
-26 if $\overline{s}_{i}<\underline{e}_{i}$ then
-27    while $tpf.time<\underline{e}_{i}$ do
-[…]
-29       $\underline{c}^{\prime}\leftarrow\underline{C}-(tpf.\overline{P}-\max(\overline{c}_{i},0))$
-30       $\overline{c}^{\prime}\leftarrow\overline{C}-(tpf.\underline{P}-\min(\underline{c}_{i},0))$
-31       if $i\in R$ then
-32          $\underline{c}^{\prime}\leftarrow\underline{c}^{\prime}+\min(\overline{c}_{i},0)$
-33          $\overline{c}^{\prime}\leftarrow\overline{c}^{\prime}+\max(\underline{c}_{i},0)$
-34
-35       $\underline{c}_{i}\leftarrow\max(\underline{c}^{\prime},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{\prime},\overline{c}_{i})$
-36       $tpf\leftarrow tpf.next$
-37
-38 else
-39    $\overline{d}^{*}\leftarrow 0$; $s^{\prime}\leftarrow\underline{s}_{i}$
-40    $\underline{c}^{*}\leftarrow\underline{C}-tpf.prev.\overline{P}+\max(\overline{c}_{i},0)$
-41    $\overline{c}^{*}\leftarrow\overline{C}-tpf.prev.\underline{P}+\min(\underline{c}_{i},0)$
-42    while $tpf.time<\overline{s}_{i}$ do
-43       $\overline{d}^{*}\leftarrow\max(tpf.time-s^{\prime},\overline{d}^{*})$
-44       if $tpf.\underline{P}+\max(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tpf.\overline{P}+\min(\overline{c}_{i},0)<\underline{C}$ then
-45          $s^{\prime}\leftarrow tpf.next.time$
-46
-47       else
-[…]
-49
-50       $\underline{c}^{*}\leftarrow\min(\underline{C}-tpf.\overline{P}+\max(\overline{c}_{i},0),\underline{c}^{*})$
-51       $\overline{c}^{*}\leftarrow\max(\overline{C}-tpf.\underline{P}+\min(\underline{c}_{i},0),\overline{c}^{*})$
-52       $tpf\leftarrow tpf.next$
-53
-54    $\overline{d}^{*}\leftarrow\max(\overline{e}_{i}-s^{\prime},\overline{d}^{*})$; $\overline{d}_{i}\leftarrow\min(\overline{d}^{*},\overline{d}_{i})$
-55    $\underline{c}_{i}\leftarrow\max(\underline{c}^{*},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{*},\overline{c}_{i})$
-56
-[…]
-Each task is processed in three steps:
-First, the Forbid rule is used to adjust the minimum start time of the task by iterating forward over the time points in the interval $[\underline{s}_{i},\min(\overline{s}_{i},\underline{e}_{i}))$ in the loop at Line 3.
-Second, the Forbid rule is used to adjust the maximum end time of the task by iterating backward over the interval $[\max(\overline{s}_{i},\underline{e}_{i}),\overline{e}_{i})$ in the loop at Line 3.
-The last step iterates over the time points of the remaining part of the time window.
-It depends if the task has a fixed part or not.
-In the former case, the loop at Line 3 iterates over the time points in the fixed part of the task and adjusts its height following the 1st case of the Height rule.
-In the case where the task has no fixed part, the loop at Line 3 is executed.
-The height of the task is adjusted according to the 2nd case of the Height rule and its maximum length is adjusted following the Length rule.
+    $\overline{d}^{*}\leftarrow\max(\overline{e}_{i}-s^{\prime},\overline{d}^{*})$; $\overline{d}_{i}\leftarrow\min(\overline{d}^{*},\overline{d}_{i})$
+    // Apply Length rule
+    $\underline{c}_{i}\leftarrow\max(\underline{c}^{*},\underline{c}_{i})$; $\overline{c}_{i}\leftarrow\min(\overline{c}^{*},\overline{c}_{i})$
+    // Restrict min/max task consumption
+26
+Algorithm 3 timetabling($T,\underline{C},\overline{C}$)
 In order to compute the maximum length of the task, we maintain the variables $\overline{d}^{*}$ which contains the maximum length found so far and $s^{\prime}$ which corresponds to the earliest time at which the task can start and span without capacity violation until the current time point.
@@
 Similarly, the minimum and maximum available heights are computed as $\underline{c}^{*}$ and $\overline{c}^{*}$ and used to update the height of the task if it has no fixed part (Line 3).
-Note that these variables are initialized based on the height available at the time point before the the start of Loop 3 in order to consider the whole minimum overlapping interval of the task.
+Note that these variables are initialized based on the height available at the time point before the start of Loop 3 in order to consider the whole minimum overlapping interval of the task.
 The Mandatory rule is enforced over the whole time window of each task.
@@
 It is called in each loop, at Lines 3, 3, 3 and 3.
-1 if  #$\#fp>0\wedge(tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C})$  then
-2 $p_{i}\leftarrow true$
-3 $\overline{s}_{i}\leftarrow\min(tp.time,\overline{s}_{i})$
-4 $\underline{e}_{i}\leftarrow\max(tp.next.time,\underline{e}_{i})$
-5 $deficit\leftarrow\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
-6 $overload\leftarrow\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
-7 if $deficit>0$ then $\underline{c}_{i}\leftarrow\max(deficit,\underline{c}_{i})$
-8 if $overload<0$ then $\overline{c}_{i}\leftarrow\min(overload,\overline{c}_{i})$
+1if $\mathit{\#fp}>0\wedge(tp.\underline{P}-\min(\underline{c}_{i},0)>\overline{C}\vee\allowbreak tp.\overline{P}-\max(\overline{c}_{i},0)<\underline{C})$ then
+2 $p_{i}\leftarrow\mathit{true}$
+3 $\overline{s}_{i}\leftarrow\min(tp.\mathit{time},\overline{s}_{i})$
+4 $\underline{e}_{i}\leftarrow\max(tp.\mathit{next}.\mathit{time},\underline{e}_{i})$
+5 $\mathit{deficit}\leftarrow\underline{C}-(tp.\overline{P}-\max(\overline{c}_{i},0))$
+6 $\mathit{overload}\leftarrow\overline{C}-(tp.\underline{P}-\min(\underline{c}_{i},0))$
+7 if $\mathit{deficit}>0$ then $\underline{c}_{i}\leftarrow\max(\mathit{deficit},\underline{c}_{i})$
+8 if $\mathit{overload}<0$ then $\overline{c}_{i}\leftarrow\min(\mathit{overload},\overline{c}_{i})$
 9
-[…]
+Algorithm 4 checkIfMandatory($i,tp,\underline{C},\overline{C}$)
 The worst-case time complexity of Algorithm 3 is $\mathcal{O}(n^{2})$.
@@
 Thus, such tasks do not need to be processed which reduces the number of time points in the profile.
-To do so, we use the same process as the Fruitless Fixed Tasks Removal of (Gay, Hartert, and Schaus 2015a).
+To do so, we use the same process as the Fruitless Fixed Tasks Removal of (GayHS15).
 Example 0.
 When executing the timetabling algorithm on the same tasks as in Example 1, the adjustments are:
-  1. 1.
-    When processing task $A$ at time point 111, its maximum height $\overline{c}_{A}$ is adjusted to 111 by the Height rule.
-  2. 2.
-    When processing task $B$ at time point 222, the task is detected as non-feasible at this time point by the Forbid rule.
-Its minimum start $\underline{s}_{B}$ is thus pushed to the next time point at time 333 where the adjustment stops as the task is feasible at this time.
+  1. (1)
+    When processing task $A$ at time point $1$, its maximum height $\overline{c}_{A}$ is adjusted to $1$ by the Height rule.
+  2. (2)
+    When processing task $B$ at time point $2$, the task is detected as non-feasible at this time point by the Forbid rule.
+Its minimum start $\underline{s}_{B}$ is thus pushed to the next time point at time $3$ where the adjustment stops as the task is feasible at this time.
 This also updates its minimum end to $\underline{e}_{B}=6$ due to the internal relation $\underline{e}_{B}\geq\underline{s}_{B}+\underline{d}_{B}$.
-  3. 3.
-    When processing task $C$ at time point 444, the task is detected as mandatory.
-The following adjustments are done: 1) The task is set to present.
-2) Its maximum starting time is set to $\overline{s}_{C}=4$ which also updates its maximum ending time to $\overline{e}_{C}=7$ due to the internal relation $\overline{e}_{C}\leq\overline{s}_{C}+\overline{d}_{C}$.
-3) Its minimum ending time is set to $\underline{e}_{C}=5$.
-4) As the value $overload=-1$ is negative, the maximum height of the task is set to $\overline{c}_{C}=-1$.
+  3. (3)
+    When processing task $C$ at time point $4$, the task is detected as mandatory. The following adjustments are done:
+    1. (a)
+      The task is set to present.
+    2. (b)
+      Its maximum starting time is set to $\overline{s}_{C}=4$ which also updates its maximum ending time to $\overline{e}_{C}=7$ due to the internal relation $\overline{e}_{C}\leq\overline{s}_{C}+\overline{d}_{C}$.
+    3. (c)
+      Its minimum ending time is set to $\underline{e}_{C}=5$.
+    4. (d)
+      As the value overload=1overload=-1 is negative, the maximum height of the task is set to $\overline{c}_{C}=-1$.
     Note that even if task $B$ is processed before task $C$ and its fixed part increased, task $C$ is adjusted based on the profile range which has been computed based on the state of task $B$ at the start of the propagation.
 The state of the tasks after timetabling is:
-$A$
-: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=1,\allowbreak p=true\rangle$
-$B$
-: $\langle s=[3,4],\allowbreak d=[3,4],\allowbreak e=[6,7],\allowbreak c=2,\allowbreak p=true\rangle$
-$C$
-: $\langle s=[3,4],\allowbreak d=[1,3],\allowbreak e=[5,7],\allowbreak c=[-2,-1],\allowbreak p=true\rangle$
-Related Work
-In (Beldiceanu and Carlsson 2002), a timetabling filtering for the GeneralizedCumulative constraint with alternative resources was introduced.
-It collects a list of tasks at relevant time-point before filtering them.
-Our method is more in line with the strategy described in (Gay, Hartert, and Schaus 2015a) for the classical cumulative, which also constructs a profile and then confronts each task with it.
-We instead use the Profile data structure (Gingras and Quimper 2016), to prune the tasks forward and backward.
-This reduces the number of calls required to reach the fix-point of the constraint.
-Contrary to (Gay, Hartert, and Schaus 2015a), which uses just a list of rectangles for the profile, our algorithm finds the starting rectangle in $\mathcal{O}(1)$ thanks to the Profile.
-Consequently, despite a worst-case quadratic time complexity per call similar to that described in (Beldiceanu and Carlsson 2002) and (Gay, Hartert, and Schaus 2015a), the proposed algorithm can be faster.
-Furthermore, our algorithm also prunes the height and the maximum length of tasks.
-More details on the filtering differences with the other implementations from the literature are provided in the technical Appendix.
-Experimental Results
-Our approach and filtering algorithm was implemented in Java on MaxiCP, an extension of the MiniCP (Michel, Schaus, and Van Hentenryck 2021) solver.
-We compare it with the closed-source state-of-the-art commercial solver CPOptimizer (Laborie et al.
-2018) as well as the open-source solver Gecode (Schulte, Lagerkvist, and Tack 2006) that implements a version of the GeneralizedCumulative constraint of (Beldiceanu and Carlsson 2002).
+AA:
+: $\langle s=[0,1],\allowbreak d=[3,4],\allowbreak e=[3,4],\allowbreak c=1,\allowbreak p=\mathit{true}\rangle$
+$B$:
+: $\langle s=[3,4],\allowbreak d=[3,4],\allowbreak e=[6,7],\allowbreak c=2,\allowbreak p=\mathit{true}\rangle$
+$C$:
+: $\langle s=[3,4],\allowbreak d=[1,3],\allowbreak e=[5,7],\allowbreak c=[-2,-1],\allowbreak p=\mathit{true}\rangle$
+Figure 5 shows the state of the tasks as well as the adjustments done.
+Figure 5. Representation of the tasks of Example 1 after timetabling. Adjustments are shown as gray arrows when applicable.
+4. Related Work
+Some forms of the generalized cumulative constraint are available in both open-source and commercial constraint programming systems.
+We distinguish between those that use a multi-resource API, based on the work of (BeldiceanuC02), and those that offer a conditional task interval API with cumulative functions.
+4.1. Generalized multi-resource constraint
+In (BeldiceanuC02), a timetabling filtering algorithm for the GeneralizedCumulative constraint with $k$ alternative resources was introduced.
+This constraint does not rely on conditional task intervals or cumulative functions (which were only introduced later in (LaborieRSV09)); instead, it operates directly on arrays of integer variables.
+The signature of the constraint is:
+$cumulatives(s[1,n],d[1,n],e[1,n],c[1,n],m[1,n],\underline{C}[1,k],\overline{C}[1,k])$
+where $s$, $d$, $e$, and $c$ denote respectively the start times, durations, end times, and consumptions (which can be positive or negative).
+The variables $m$ indicate, for each task, the resource on which it executes, with domain on $\{1,\ldots,k\}$.
+Finally, $\underline{C}[j],\overline{C}[j]]$ specifies the capacity range of resource $j$.
+Although this constraint does not directly support conditional task intervals, it is straightforward to connect the two modeling approaches.
+Indeed, a dummy resource can be used to indicate if a task is not present in the multi-resource formulation.
+Conversely, the conditional task approach supports multiple resources by using an additional alternative constraint (LaborieR08) that represent the choice of alternative resources.
+This requires duplicating the activity once for each alternative resource, but it has the advantage that each alternative task can be filtered independently during propagation.
+In contrast, the multi-resource formulation propagation can only exploit the minimum over the different possible starts across all resources, which can lead to a weaker filtering.
+The filtering algorithm of (BeldiceanuC02) proceeds by collecting a set of tasks at relevant time points before performing propagation.
+This algorithm is implemented in the open-source solver Gecode (schulte2006gecode) and according to its documentation, the same algorithm is also implemented in Sicstus (sicstusmanual).
+The algorithm presents some substantial differences with our algorithm.
+First, our approach is closer to the strategy described in (GayHS15) for the classical cumulative constraint, which constructs a profile and then tests each task against it.
+We rely on the Profile data structure (GingrasQ16) to prune the tasks forward and backward, thus reducing the number of propagation calls required to reach a fixpoint.
+In contrast to (GayHS15), which represents the profile as a simple list of rectangles, our algorithm can retrieve the starting rectangle in $\mathcal{O}(1)$ thanks to the Profile.
+Consequently, although the worst-case quadratic time complexity per call remains similar to that of (BeldiceanuC02) and (GayHS15), our algorithm can be faster in practice.
+Second, our algorithm can possibly prune more the height and the duration attributes of the tasks.
+These filtering differences are detailed next.
+Backward propagation
+For the adjustment of the maximum end time ($\overline{e}$) of tasks due to the Forbid rule, when a task requires two or more consecutive adjustments, our algorithm performs all the adjustments in a single call to the timetabling algorithm compared to the version of (BeldiceanuC02) where only one adjustment is done per call.
+Indeed, in (BeldiceanuC02), detection and adjustment are performed only in a forward fashion.
+Thus, the filtering of the maximum end time of tasks is done only in regards to the current end of the task time window.
+This means that if a task has its maximum end time adjusted, whether or not the maximum end time can be further adjusted will be considered in the next call to the filtering algorithm.
+In practice, this may lead to situations where the filtering algorithm of (BeldiceanuC02) needs several consecutive calls in order to completely adjust the end time of a task and reach a fixed point.
+In contrast the algorithm presented in this paper does such adjustments in a single pass as illustrated in Example 1.
+Example 0.
+Let us consider three tasks:
+  • Task A is fixed and present, starts at 4, ends at 5 and has a height of 1.
+  • Task B is fixed and present, starts at 7, ends at 8 and has a height of 1.
+  • Task C is present, has a minimum start of 0, a maximum end of 10, a fixed length of 3 and a height of 1.
+The maximum capacity of the resource $\overline{C}$ is 1.
+Figure 6 shows the three tasks before propagation.
+The dashed blue rectangle represents the time window $\underline{s},\overline{e})$ of task C.
+Figure 6. Tasks of example 1
+During the backward propagation of our timetabling algorithm, task C is detected as infeasible at times 8 then 5, and its maximum end time is successively updated to 7 then 4 in a single call.
+The propagation from (BeldiceanuC02) iterates the profile in one direction and detects the infeasibility of task C only at time 8 and adjusts its maximum end at 7 during the first call.
+A second call to the algorithm is necessary in order to adjust the maximum end at time 4.
+Height Adjustment Difference
+For height adjustment, the difference occurs in the case where a task has no fixed part (second case of the Height filtering rule).
+In this case, our algorithm considers the maximum height available over the minimum overlapping interval of the task (gay2015time).
+The propagation from (BeldiceanuC02) uses another height adjustment rule (see Alg.
+4 in their paper).
+It considers the maximum available height only if the task examined has both its earliest completion time ($\underline{e}$) and its latest start time ($\overline{s}$) in the same event interval.
+That means that if one or more other tasks affect the available height between the $\underline{e}$ and $\overline{s}$ of a task, leading them to occur during different events, the height of the task is not adjusted.
+Example 2 illustrates this.
+Example 0.
+Let us consider three tasks:
+  • Task A is fixed and present, starts at 4, ends at 12 and has a height of 2.
+  • Task B is fixed and present, starts at 6, ends at 10 and has a height of -1.
+  • Task C has a minimum start of 0, a maximum end of 16 and a fixed length of 6. Its height is in the interval $[1,4]$
+The maximum capacity of the resource $\overline{C}$ is 4.
+Figure 7 shows the three tasks before propagation.
+The dashed blue rectangle represents task C maximum energy.
+The dotted black line shows the minimum profile $\underline{P}$.
+Figure 7. Tasks of example 2
+During the execution of the timetabling algorithm, the maximum height of task C is adjusted to 3 as this value corresponds to the maximum available height over its minimum overlapping interval (from time 5 to time 10).
+This filtering does not occur with the GeneralizedCumulative constraint from (BeldiceanuC02) as the $\underline{e}_{C}$ and $\overline{s}_{C}$ of task C are not in the same event due to the presence of task B.
+If the task B is removed, then both algorithms obtain the same filtering as the height of task C is adjusted to 2.
+Length Adjustment Difference
+The difference in filtering on the maximum length adjustment of the tasks occurs if the profile prevents a task to be present in at least two different parts of its time window.
+In this case, the algorithm presented in this paper identifies the longest interval in the profile where the task can fit and adjusts the task’s maximum length accordingly.
+In contrast, the filtering of (BeldiceanuC02) one single conflict point at a same time.
+When encountering conflicting time point in the profile, the task maximum length is updated based on the maximum length between the conflicting time point and either the minimum start or the maximum end of the task.
+Example 3 illustrates this.
+Example 0.
+Let us consider three tasks:
+  • Task A is fixed and present, starts at 3, ends at 6 and has a height of 3.
+  • Task B is fixed and present, starts at 10, ends at 14 and has a height of 3.
+  • Task C has a minimum start of 0, a maximum end of 16 and a height of 2. Its length is in the interval $[2,16]$
+The maximum capacity of the resource $\overline{C}$ is 4.
+Figure 8 shows the tree tasks before propagation.
+The dashed blue rectangle represents task C maximum energy.
+The black dotted line shows the maximum capacity ($\overline{C}$).
+In this example, both tasks A and B prevent task C to be present at the same time.
+Figure 8. Tasks of example 3
+During the execution of the timetabling algorithm, the maximum length of task C is adjusted to 4 as it is the length of the longest interval where the task can be placed (from time 6 to time 10).
+When filtering with the propagation algorithm from (BeldiceanuC02), the length is instead adjusted to 10, since this corresponds to both the length available after task A and the length available before task B.
+4.2. Closed source solvers
+Both CPOptimizer (laborie2018ibm) and OptalCP (OptalCP) have an implementation of the cumulative functions modeling paradigm but the details of their filtering algorithms has not been published and both solvers are closed-source.
+Therefore, the exact filtering used for their range constraints on cumulative functions is difficult to assert.
+Based on the filtering and the number of backtracks on some examples, we observed that CPOptimizer achieves a filtering similar to the Gecode implementation when a cumulative function requires a generalized cumulative constraint.
+For specific cases, like when a cumulative function models a non-overlap constraint, the constraint is reformulated by CPOptimizer into a standard disjunctive constraint for which strong dedicated filtering exists (such as edge-finding, not-first/not-last, etc.).
+Contrary to ours, the CPOptimizer modeling API does not allow cumulative functions with a negative height.
+Indeed, while cumulative tasks can have a negative contribution, a cumulative function with a negative height will cause a failure in CPOptimizer.
+This is equivalent to having a constraint $f\geq 0$ on any cumulative function $f$ used in the model.
+It is not known whether this limitation is due to a design choice or a limitation in the filtering algorithm used for the constraint.
+5. Experiments
+Our approach and filtering algorithm are open source and available in MaxiCP (MaxiCP2024), a Java solver that extends MiniCP (MiniCP).
+We compare it with the closed-source state-of-the-art commercial solver CPOptimizer (laborie2018ibm) as well as the open-source solver Gecode (schulte2006gecode) that implements a version of the GeneralizedCumulative constraint of (BeldiceanuC02).
 As cumulative functions are not available in Gecode, we carefully modeled them in a way similar to how they would be flattened by Algorithm 1.
@@
 The experiments were carried out on a MacBook Pro M3 with 32GB of memory.
-A description of the models is available in the technical appendix, and the sourcecodes are available in the supplementary material.
-Figure 4 shows the cumulative number of instances solved within a virtual time limit for the three problems.
-The appendix also presents a pairwise comparison of solvers on the number of backtracks for instances that were commonly solved.
-RCPSP-CPR
-The Resource Constrained Project Scheduling Problem with Consumption and Production of Resources (RCPSP-CPR) is a variant of the RCPSP with additional reservoir resources that must always be kept positive (Koné et al.
-2013).
+The source code of the models is available in the supplementary material.
+For each problem, we report the results as a plot of the cumulative number of instances solved within a virtual time limit.
+We also provide a pairwise comparison of solvers on the number of backtracks for instances that were commonly solved.
+Each model involves a set of conditional task intervals $i\in T$.
+Recall that the attributes for each interval are $s_{i}$ the start time, $d_{i}$ the duration, $e_{i}$ the end time, $p_{i}$ the status of a task (present or absent) when applicable.
+5.1. The RCPSP with Consumption and Production of Resources (RCPSP-CPR)
+The Resource-Constrained Project Scheduling Problem with Consumption and Production of Resources (RCPSP-CPR) (kone2013comparison) is an extension of the classic Resource-Constrained Project Scheduling Problem (RCPSP) (Dike_RCPSP; DING_2023_RCPSP) with additional reservoir resources that must always be kept positive.
 Each task is present and, in addition to classical renewable resource consumption (modeled with pulse functions), tasks also consume a specified amount of reservoir resources at their start and produce a specified amount at their end (modeled with stepAtStart and stepAtEnd).
 The objective is to minimize the makespan.
-We experiment on 55 instances with 15 tasks adapted from (Koné et al.
-2013).
+Model
+$T$ denotes the set of tasks, $Rn$ is the set of renewable resources, $Rs$ is the set of reservoir resources, and $P$ is the set of precedences between pairs of tasks.
+The capacity of each renewable resource $k\in Rn$ is denoted $C_{k}$ and the initial level of each reservoir resource $u\in Rs$ is denoted $C^{r}_{u}$.
+Each task $i$ consumes $c_{ki}$ unit of the renewable resource $k\in Rn$, consumes $c^{-}_{ui}$ unit of reservoir resource $u\in Rs$ at its start time and produces $c^{+}_{ui}$ unit of reservoir resource $u\in Rs$ at its end time.
+The model is written as:
+(11) $\mathit{minimize}\;\max\limits_{i\in T}e_{i}$
+subject to
+(12) $e_{i}\leq s_{j}$ $\forall(i,j)\in P$
+(13) $ren_{k}=\sum\limits_{i\in T}pulse(i,c_{ki})$ $\forall k\in Rn$
+[…]
+(15) $ren_{k}\leq C_{k}$ $\forall k\in Rn$
+(16) $res_{u}\geq 0$ $\forall u\in Rs$
+The objective (11) is to minimize the makespan.
+The constraints (12) ensure the precedences between tasks.
+The renewable and reservoir resources are constrained in (13) and (14) respectively.
+The resource constraints are posted at (15) and (16) respectively.
+The search consists in a static binary branching that selects variables in the order in which tasks are declared in the instance file.
+The first un-assigned start variables is selected and assigned to its minimum value in the left branch.
+The right branch removes the minimum value from the domain.
+Results
+We experiment on 55 instances with 15 tasks adapted from (kone2013comparison).
 A time limit of 600s is set per instance for finding and proving optimality.
 As shown in Figure 9, MaxiCP solves more instances than Gecode and roughly the same set of instances as CPOptimizer.
-A backtrack analysis between CPOptimizer and MaxiCP reveals that the number of backtracks is nearly identical, suggesting that CPOptimizer is simply faster on this problem, as both solvers achieve similar levels of filtering on this problem.
-SMIC
-The Single Machine with Inventory Constraints (SMIC) problem was introduced in (Davari et al.
-2020).
-Each task has a release date, a fixed duration, and a positive or negative inventory consumption that occurs at the start of the activity (modeled with stepAtStart).
-The inventory starts at a given value and must stay within a specified range.
+Figure 9. Number of solved instances in function of time for the RCPSP-CPR Problem.
+Figure 10 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks.
+The comparison between CPOptimizer and MaxiCP reveals that the number of backtracks is nearly identical, suggesting that CPOptimizer is simply faster on this problem, as both solvers achieve similar levels of filtering on this problem, but MaxiCP requires less backtracks than Gecode form some of the instances.
+Figure 10. Failure comparison for the RCPSP-CPR
+5.2. The Single Machine with Inventory Constraints (SMIC)
+The Single Machine with Inventory Constraints (SMIC) problem was introduced in (smic).
+Each task has a release date $release_{i}$, a fixed duration, and a positive or negative inventory consumption $invent_{i}$ that occurs at the start of the activity (modeled with stepAtStart).
+The inventory starts at a given value $\mathit{initInventory}$ and must stay within a specified range ($0,\mathit{capaInventory}]$).
 Activities cannot overlap in time.
 The objective is to minimize the makespan.
-We experiment on instances with 10 tasks from (Davari et al.
-2020).
-The non-overlap constraint was enforced using the generalized cumulative constraint in MaxiCP (modeled using pulse) and Gecode.
+Model
+The model is written as:
+(17) $\mathit{minimize}\;\max_{i\in T}e_{i}$
+subject to
+(18) $s_{i}\geq\mathit{release_{i}},$ $\forall i\in T$
+(19) $res=\mathit{step}(0,\mathit{initInventory})+\sum_{i\in T}\mathit{stepAtStart}(i,invent_{i}),$ $\forall i\in T$
+(20) $0\leq res\leq\mathit{capaInventory}$
+(21) $\mathit{noOverlap}=\sum_{i\in T}pulse(i,1),$ $\forall i\in T$
+(22) $\mathit{noOverlap}\leq 1$
+The objective (17) is to minimize the makespan (i.e., the time at which all tasks are completed).
+The release dates are ensured by the relation (18).
+The reservoir resource is constrained in relations (19) and (20) to ensure that the capacity of the inventory is not violated.
+The noOverlap constraint is enforced with the relations (21) and (22) to ensure the no-overlapping of activities.
+Notice that those are modeled using cumulative functions rather than with global dedicated non-overlap constraints on task intervals to ensure that the results are solely impacted by the filtering of the cumulative constraint rather than by the possibly different filtering of the non-overlap constraint.
+This, unfortunately, did not appear possible for CPOptimizer which recognizes that the cumulative function is used to model a non-overlap constraint and automatically reformulates it using a non-overlap constraint with stronger filtering.
+The search consists in a static branching heuristic that selects variables in the order in which tasks are declared in the instance file.
+The first un-assigned start variables is selected and assigned to its minimum value in the left branch.
+The right branch removes the minimum value from the domain.
+Results
+We experiment on instances with 10 tasks from (smic).
 A time limit of 600s is set per instance for finding and proving optimality.
-As shown in Figure 4 CPOptimizer solves more instances.
-A backtrack analysis showed that CPOptimizer requires significantly fewer backtracks on this problem, while Gecode and MaxiCP are not on par for this problem in terms of time and number of backtracks.
-Unfortunately the results for CPOptimizer for this problem are biased as it appears to reformulate the cumulative function for modeling the non-overlap into a standard disjunctive constraint for which strong dedicated filtering exist such as (Vil´ım 2004).
-MESP
-The Maximum Energy Scheduling Problem (MESP) involves optional tasks with variable durations and resource demands, which can be positive or negative, on a single resource with fixed capacity.
+As shown in Figure 11 CPOptimizer solves more instances while MaxiCP and Gecode roughly have the same behavior on this problem.
+Figure 11. Number of solved instances in function of time for the SMIC Problem
+Figure 12 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks.
+This analysis shows that CPOptimizer requires significantly fewer backtracks on this problem, while Gecode and MaxiCP are not on par for this problem in terms of time and number of backtracks.
+Unfortunately the results for CPOptimizer for this problem are biased as it appears to reformulate the cumulative function for modeling the non-overlap into a standard disjunctive constraint for which strong dedicated filtering exist such as (vilim2004nlog).
+Figure 12. Failure comparison for the SMIC
+5.3. The Maximum Energy Scheduling Problem (MESP)
+The Maximum Energy Scheduling Problem (MESP) consists in scheduling optional tasks with variable durations and resource demands, which can be positive or negative, on a single resource with fixed capacity $\overline{C}$.
+The energy of a task is the product of the duration and the demand.
 The objective is to maximize the total positive energy consumption of the resources while ensuring that the resource capacity is not exceeded at any time.
 All filtering rules are beneficial for this problem, making it the most relevant benchmark for evaluating the proposed filtering algorithm.
+The model is written as:
+(23) $\mathit{maximize}\;\sum\limits_{i\in T\mid p_{i}=\mathit{true}}\max(c_{i},0)\cdot d_{i}$
+subject to
+(24) $\mathit{resource}=\sum\limits_{i\in T}\mathit{pulse}(i,c_{i})$
+(25) $\mathit{resource}\leq\overline{C}$
+The objective (23) is to maximize the sum of the positive energies of the tasks that are present. The constraint (24) sets up the reservoir resource and its capacity is constrained by (25).
+Note that the resource demand variables of tasks $c_{i}$ may have negative values in their domain in this problem.
+As explained in Filtering Differences section, CPOptimizer does not allow cumulative functions to have a negative height.
+In order to accommodate this restriction, the model for CPOptimizer has been adapted in the following way: The cumulative function resourceresource is shifted up in height at the time 0 by a step function of height $neg^{*}=\sum_{i\in T}-\min(\underline{c}_{i},0)$ which corresponds to the sum of all maximum negative values of the tasks resources demands.
+The capacity $\overline{C}$ of the resource is also shifted by the same value.
+In formal terms, constraint (24) is changed to: $resource=step(0,neg^{*})+\sum\limits_{i\in T}pulse(i,c_{i})$ and constraint (25) is changed to $resource\leq\overline{C}+neg^{*}$.
+This ensures that the cumul function resourceresource remains non negative, as required by CPOptimizer.
+The search heuristic consists in selecting tasks in the fixed order of their declaration in the instance file. For each task, the following variables are fixed in this order:
+  • Its presence $p_{i}$ to true
+  • Its demand $c_{i}$ to the maximum value in the domain
+  • Its duration $d_{i}$ to the maximum value in the domain
+  • Its end time $e_{i}$ to the maximum value in the domain
+Two branches are created: The left branch fixes the value while the right branch removes it from the domain. The search stops at the first solution found.
+Results
 We generated 60 instances with a number of tasks ranging between 6 and 12800.
+A time limit of 2,000s and a memory limit of 16GB is set per instance for finding a feasible solution.
 Given the large number of tasks in some instances, this experiment is designed to test the scalability of the filtering.
-A time limit of 2,000s and a memory limit of 16GB is set per instance for finding a feasible solution.
 As can be observed in Figure 13, our approach performs very well with more instances solved.
 Both Gecode and CP Optimzer fail to solve largest instances due to timeouts and memory limits respectively.
-A backtrack analysis shows that our approach avoids backtracking, unlike the other solvers, and that CP Optimizer requires fewer backtracks than Gecode although it solves less instances.
-Figure 4: Number of solved instances in function of time for the RCPSP-CPR, SMIC and MESP Problems.
-Conclusion
-We explained how the GeneralizedCumulative constraint can be used to implement modeling paradigm of cumulative functions for scheduling problems.
-We also introduced a time-tabling filtering algorithm for the GeneralizedCumulative constraint.
-The experimental results show that the proposed approach is scalable and competitive with other solvers on three cumulative scheduling problems with varying characteristics.
+Figure 13. Number of solved instances in function of time for the MESP Problem.
+Figure 14 reports for each pair of solvers, for each instance solved commonly by both solvers, an $x,y$- plot of the number of backtracks.
+As can be seen, our approach avoids backtracking, unlike the other solvers.
+Interestingly, CPOptimizer requires fewer backtracks than Gecode although it solves less instances.
+Figure 14. Failure comparison for the MESP
+5.4. Why not compare with solvers using Minizinc?
+The MiniZinc modeling language provides a cumulatives constraint that allows for negative consumption.
+Unfortunately, with the exception of Sicstus (CarlssonM12) and Gecode (schulte2006gecode), this constraint is decomposed in other solvers.
+Also the MiniZinc API for this constraint does not support explicit access to the end variables.
+Being able to access the end variables directly rather than relying on the relation start + duration = end is crucial for the problems considered, as the end variables can be constrained independently, as explained in the next example.
+Example 0.
+Assume a task with the following properties: $s=[3,7]$, $d=[1,5]$, $e=[8,8]$.
+The mandatory part of this task is $\overline{s},\underline{e}-1]=[7,7]$.
+If the end variable is not available, the mandatory part inferred from $s$ and $d$ is $\overline{s},\underline{s}+\underline{d}-1]=\phi$.
+As a result, the propagation when using MiniZinc for these problems can be much weaker, making the comparison unfair.
+This was confirmed experimentally and can be verified with the provided MiniZinc models in the appendix.
+6. Conclusion
+We presented an implementation of the cumulative functions modeling paradigm for scheduling problems.
+Our implementation applies a flattening procedure to cumulative function expressions, producing a set of activities with positive or negative resource consumption, which are then passed to a GeneralizedCumulative constraint.
+We introduced a novel time-tabling filtering algorithm for the GeneralizedCumulative constraint that operates on conditional task intervals and supports negative resource consumption.
+This algorithm is simpler and achieves stronger pruning than the one proposed by (BeldiceanuC02).
+Experimental results demonstrate that the proposed approach is scalable and competitive with other solvers on three cumulative scheduling problems with diverse characteristics, including optional tasks and negative resource consumption.
+As future work, we plan to extend the filtering algorithm to generalize the Overload/Underload checking and edge-finding rules in the context of producers and consumers.
+We also intend to explore explanations for the generalized cumulative constraint, similar to those proposed in (schutt2011explaining), and to investigate whether a CP-SAT solver such as OR-Tools (perron_et_al) could benefit from them.